### v1

# Hydrodynamic noise in one dimension: projected Kubo formula and its vanishing in integrable models

Hydrodynamic noise in one dimension:
projected Kubo formula
and its vanishing in integrable models
Benjamin Doyon
Department of Mathematics, King’s College London, Strand, London WC2R 2LS, U.K.
Hydrodynamic noise is the Gaussian process that emerges at larges scales of space and time in many-body systems. It arises by the central limit theorem applied to local microcanonical averages, representing the degrees of freedom that have been forgotten when projecting coarse-grained observables onto conserved quantities. It comes with “bare” diffusion terms. In one dimension of space, nonlinearities of the hydrodynamic equation are relevant (from a renormalisation perspective), usually giving rise to hydrodynamic superdiffusion. But in linearly degenerate systems, where the relevant nonlinearity vanishes, the diffusive scaling stays intact. Nevertheless, anomalies remain. We show that in such systems, the noise covariance is determined in terms of a modification of the Kubo formula, where effects of ballistic long-range correlations have been subtracted. This is the projected Onsager matrix, in which so-called quadratic charges are projected out. We show that the Einstein relation holds, giving a projected bare diffusion, and that the remaining nonlinearities are tamed by a point-splitting regularisation. Putting these ingredients together, we obtain an exact and well-defined hydrodynamic fluctuation theory in the ballistic scaling of space-time, for the asymptotic expansion in the inverse variation scale, including the first subleading (diffusive-scale) corrections beyond large deviations. This is expressed as a stochastic PDE. We then obtain the anomalous hydrodynamic equation, which takes into account separately long-range correlations and bare diffusion. Using these result, in integrable systems, we show that hydrodynamic noise must be absent, as was conjectured recently.
1 Introduction
At large scales of space and time, degrees of freedom in many-body systems separate into “fast” and “slow” ones. Fast ones can be assumed to quickly relax throughout space-time, at each time taking values that are determined by the spatial profile of slow ones. This is the projection principle of Mori and Zwanzig [1, 2], and slow degrees of freedom are associated with conserved quantities [3, 4, 5, 6]. At the linearised level, it is called the Boltzmann-Gibbs principle, and says that fluctuation fields for arbitrary observables linearly project onto fluctuation fields for conserved densities [6].
More precisely, on short times we may assume coarse-grained conserved densities to be fixed, while generic coarse-grained observables fluctuate a lot, thus converging to their microcanonical values, functions of the conserved densities, in each “fluid cell”. This is exact in the strict limit where spatial variations (variations of particle or energy densities, etc.) are on large wavelengths $\ell$, as fluid cells can be taken as large as we want. But looking at correction terms in the hydrodynamic expansion in $1/\ell$, this projection is not perfect. The microcanonical averages on fluid cells receive many corrections. In particular, by the central limit theorem, there must be an emergent Gaussian white noise associated to each coarse-grained observable. This noise is delta-correlated in space-time, but correlated amongst different observables. Its covariance is usually assumed to be given by the Kubo formula for the Onsager matrix, and by the fluctuation-dissipation theorem, this noise must be associated with dissipation, giving viscous and diffusive terms by the Einstein relation [4, 5, 6, 7]. This is the basis for the theory of fluctuating hydrodynamics. It is valid for both deterministic and stochastic systems. In particular, in the latter, the emergent noise is in general not simply reIated to the microscopic noise. In purely diffusive systems, where fluxes vanish, and under the diffusive scaling of space and time $\vec{{\rm x}}\propto\ell,\,\mathrm{t}\propto\ell^{2}$, the hydrodynamic expansion in fact starts at the diffusive level, and this is the basis for the Macroscopic Fluctuation Theory (MFT) [8], a large-deviation theory.
In one dimension of space, if fluxes are non-zero, such as in Galilean or relativistic gas or in integrable models, their nonlinearity as functions of conserved densities are generically relevant in the sense of the renormalisation group, and leads to superdiffusion. For instance, take a system with a single conservation law $\partial_{\mathrm{t}}q+\partial_{\mathrm{x}}j=0$. One writes the coarse-grained current $\overline{j}$ as a function of the coarse-grained conserved density $\overline{q}$
$\overline{j}=\mathsf{j}(\overline{q})-\frac{1}{2}\mathfrak{D}\partial_{\mathrm{x}}\overline{q}+\xi$ (1)
where $\mathsf{j}(\overline{q})$ is the microcanonical average – the flux –, and $\xi$ is the noise, $\langle\!\langle\xi(\mathrm{x},\mathrm{t})\xi(\mathrm{x}^{\prime},\mathrm{t}^{\prime})\rangle\!\rangle=\mathfrak{L}\delta(\mathrm{x}-\mathrm{x}^{\prime})\delta(\mathrm{t}-\mathrm{t}^{\prime})$, with the associated diffusion function $\mathfrak{D}=\mathfrak{L}/\chi$ where $\chi$ is the susceptibility. The fluctuating hydrodynamics is the stochastic PDE
$\partial_{\mathrm{t}}\overline{q}+\partial_{\mathrm{x}}\Big{(}\mathsf{j}(\overline{q})-\frac{1}{2}\mathfrak{D}\partial_{\mathrm{x}}\overline{q}+\xi\Big{)}=0.$ (2)
This equation is generically singular and not well-defined. However, it is known that if the second derivative $\mathsf{j}^{\prime\prime}$ is non-vanishing, then the result is in the KPZ universality class under the appropriate scaling of space and time, giving superdiffusion. With more conservation laws, $\partial_{\mathrm{t}}q_{i}+\partial_{\mathrm{x}}j_{i}=0$, $i=1,2,\ldots$, similar effects arise if certain coefficients, combinations of flux derivatives, do not vanish, see [9]. These coefficients can be written as [10]
$\frac{\partial^{2}\mathsf{j}_{i}}{\partial\overline{q_{I}}^{2}}$ (3)
where densities $q_{I}$ are in the “normal mode” basis (see Appendix A).
But there is a class of models where superdiffusion does not arise: linearly degenerate systems. Linear degeneracy means that hydrodynamic modes do not self-interact: the hydrodynamic velocity of any normal mode, does not depend on this normal mode [11, 12, 13] (see also [14]). These are still interacting, yet close to free systems. They do not develop shocks [15, 16, 17, 18], and satisfy (3) [19], amongst other properties. Integrable many-body systems have hydrodynamic equations that are linearly degenerate, and there are many other examples [20, 21, 22, 23]. Because there is no shock, the fluctuating theory at the leading order in the ballistic scaling
${\rm x}=\ell x,\quad\mathrm{t}=\ell t,\quad\ell\to\infty$ (4)
is simply the deterministic transport of initial fluctuations via the Euler equation. This is the Ballistic Macroscopic Fluctuation Theory (BMFT) [24, 25], a large-deviation theory paralleling the MFT. But beyond this, for the $1/\ell$ corrections to the large-deviation behaviour, the many-component equivalent of the fluctuating hydrodynamics (2) still generically is nonlinear. Further, it was recently conjectured that in integrable systems, there is no noise, $\mathfrak{D}=\xi=0$ [20, 21, 19], and this was proposed to hold in linearly degenerate systems more generally [19].
In this paper, we propose the exact fluctuating hydrodynamic theory for linearly degenerate systems, in the ballistic scaling (4), up to, including, the first subleading $1/\ell$ order of the asymptotic expansion. The resulting stochastic PDE, Eq. (25) below, is well defined for the asymptotic expansion in $1/\ell$. Along with the nonlinear, fluctuating Boltzmann-Gibbs principle (3.1) for generic observables, it allows one to evaluate all connected correlation functions as $\ell\to\infty$, both their leading BMFT large-deviation form, worked out in [25, 26], and the first subleading order in $1/\ell$. Likewise, a field-theoretic reformulation of the stochastic PDE, such as that proposed in [25, Sec 3.3], should provide large deviations and their first subleading corrections. These will be studied in a future work.
The crucial points of this paper are as follows. (1) We show that the fluxes are subject to a “regularisation”, coming from the non-fluctuating, non-diffusive correction to the fluid-cell microcanonical average, in inverse power of the cell’s size. This regularisation, in turn, guarantees that the stochastic PDE is well defined in the $1/\ell$ expansion, despite delta-correlations both in noise and initial conditions. It is the point-splitting regularisation first conjectured in [19]. (2) We show that the noise covariance is not given by the Kubo formula for the Onsager matrix, but by a “projected” version of it. In this version, quadratic charges [10] are projected out. These represent the effects of long-range correlations [24], coming from the nonlinearity of the flux, on the diffusive scale, as worked out in [19] in the context of integrability. The fact that flux nonlinearities give rise to diffusive contributions was first proposed as “diffusion from convection” in [27] in the context of integrability, and our proof uses a calculation that is somewhat similar. (3) We show that the diffusion matrix is connected to the projected Onsager matrix by the Einstein relation. (4) We show that in integrable models, the noise for currents vanishes. This follows from our result, along with earlier calculations of the full Onsager matrix in integrable models [28, 4, 29, 30, 31]. We also derive it from first principles using the infinite family of commuting flows that many-body integrable models admit. This shows the conjecture of [20, 21, 19]. We emphasise, however, that in generic linearly degenerate systems, the noise does not vanish – this is a property of integrability.
The projected Kubo formula for the noise is also worked out, albeit in the diffusive instead of ballistic scaling, in the simultaneous paper [23] of which we have been made aware. In particular, this is evaluated in a specific model, where it is shown to agree with microscopic calculations and to be non-zero.
We emphasise that bare diffusion is different from hydrodynamic diffusion. Linearised hydrodynamics for two-point correlatoin functions is diffusive. There, however, the hydrodynamic diffusion constant is connected by the Einstein relation to the full Onsager matrix, not the projected one (see e.g. [7]). Thus, even integrable models have diffusive two-point functions. Beyond the linearised level, however, these diffusive effects desintegrate into long-range correlations [19].
From our fluctuating hydrodynamic theory, write down the hydrodynamic equation up to, including, the diffusive order $1/\ell$, see Eqs. (48), (49) below. It is a simple modification of that proposed in [19] for integrable models. Long-range corelations and bare diffusion are explicitly taken into account.
We believe our derivation gives a better understanding of the microscopic origin of hydrodynamic noise and bare diffusion, and of the subtle principles behind fluctuating hydrodynamics more generally.
The paper is organised as follows. In Section (2) we specify our general setup, describing the conserved quantities and states. In Section 3, we express our main result for the fluctuating hydrodynamic theory, and provide arguments as to how coarse-graining give rise to the special effects seen beyond the Euler scale, including the noise. In Section 4, we prove points (1)-(3) described above and write down the hydrodynamic equation. In Section 5 we show the absence of noise at the diffusive level in integrable systems. Finally, in Section 6 we conclude.
2 Conservation laws and states
Consider a dynamical system in one spatial dimension, which may be deterministic or stochastic, classical or quantum, systems with short-range, translation invariant interactions. Local observables $o({\rm x},\mathrm{t})$ are functions on phase space, random variables of stochastic processes, or operators on a Hilbert space, etc, which are supported on finite regions in microscopic units, around position ${\rm x}\in{\mathbb{R}}$ at time $\mathrm{t}\in{\mathbb{R}}$. We will write ${\rm x}=\ell x,\,\mathrm{t}=\ell t$ for the space-time point coordinates $(x,t)$ in macroscopic unit, where $\ell$ is the large macroscopic scale.
The system admits a certain number of extensive conserved quantities [^1] Integrals without domain specification are on the full space, which is ${\mathbb{R}}$ in continuous sytems or in macroscopic units, and ${\mathbb{Z}}$ in systems with discrete space and in microscopic units..
$Q_{i}=\int{\rm d}{\rm x}\,q_{i}({\rm x},\mathrm{t}),\quad\partial_{\mathrm{t}}Q_{i}=0$ (5)
where $q_{i}({\rm x},\mathrm{t})$ is the associated local density. The index $i$ lies in some index set which may be finite or not. The index may even be continuous like in integrable models where it represents the spectral space (such as velocities of quasiparticles); we still use the notation $\sum_{i}$. We assume that there are associated local currents $j_{i}({\rm x},\mathrm{t})$ giving rise to local conservation laws:
$\partial_{t}q_{i}({\rm x},\mathrm{t})+\partial_{\rm x}j_{i}({\rm x},\mathrm{t})=0.$ (6)
When discussing hydrodynamics at the diffusive order, the concept of PT symmetry plays an important role, as first discussed in [30]. We assume that there is an involution ${\mathcal{P}T}$ of the algebra of observables that reverse space and time, and that preserves all conserved densities and their currents. Further, we will only consider local observables $o$ that are real or hemitian, and that are preserved by the PT transformation. That is,
${\mathcal{P}T}(q_{i}({\rm x},\mathrm{t}))=q_{i}(-{\rm x},-\mathrm{t}),\quad{\mathcal{P}T}(j_{i}({\rm x},\mathrm{t}))=j_{i}(-{\rm x},-\mathrm{t}),\quad{\mathcal{P}T}(o({\rm x},\mathrm{t}))=o(-{\rm x},-\mathrm{t}).$ (7)
This simplifies the discussion in subtle ways (we will use such PT invariance at one crucial point). Note that in general, there may be observables $o$ with negative eigenvalue under the ${\mathcal{P}T}$ involution; however, our results do not apply to such observables.
The macrocanonical ensemble is that which maximises entropy states under the constraints of the extensive conserved quantities. Formally, it has the Gibbs form
$\langle\cdots\rangle_{\underline{\beta}}=\lim_{\mathcal{V}\to{\mathbb{R}}}\frac{\int{\rm d}\mu\,{\rm e}^{-\sum_{i}\beta^{i}Q_{i}}\cdots}{\int{\rm d}\mu\,{\rm e}^{-\sum_{i}\beta^{i}Q_{i}}}$ (8)
where $\underline{\beta}=(\beta^{1},\beta^{2},\ldots)$ are the “Lagrange parameters”, the charges $Q_{i}=\int_{\mathcal{V}}{\rm d}^{d}{\rm x}\,q_{i}(\bm{{\rm x}})$ lie on the volume $\mathcal{V}$, and the large-volume limit $\mathcal{V}\to{\mathbb{R}}$ is assumed to exist for averages of local observables. $\mu$ is a time-invariant, homogeneous prior measure, for instance in classical Hamiltonian systems with canonical coordinates $({\rm x}_{a},{\rm p}_{a})$ one chooses ${\rm d}\mu=\sum_{N=0}^{\infty}(N!)^{-1}\prod_{a=1}^{N}{\rm d}{\rm x}_{a}{\rm d}{\rm p}_{a}$ (direct sum); while in quantum systems, $\langle\cdots\rangle_{\underline{\beta}}=\frac{{\rm Tr}\,{\rm e}^{-\sum_{i}\beta^{i}Q_{i}}\cdots}{{\rm Tr}\,{\rm e}^{-\sum_{i}\beta^{i}Q_{i}}}$. We assume that $\beta^{i}$’s are away from phase transitions, so that correlations are short-range in these states, and averages of local observables are smooth in $\underline{\beta}$. Likewise, the microcanonical ensemble is denoted (by a slight abuse of notation)
$\langle\cdots\rangle_{\underline{\mathsf{q}}}=\lim_{\mathcal{V}\to{\mathbb{R}}^{d}}\frac{\int_{Q_{i}/|\mathcal{V}|\,\in\,[\mathsf{q}_{i}-\epsilon,\mathsf{q}_{i}+\epsilon]\,\forall i}{\rm d}\mu\,\cdots}{\int_{Q_{i}/|\mathcal{V}|\,\in\,[\mathsf{q}_{i}-\epsilon,\mathsf{q}_{i}+\epsilon]\,\forall i}{\rm d}\mu}$ (9)
where $\epsilon\to 0$ as $\mathcal{V}\to{\mathbb{R}}$ in an appropriate fashion [32, 33]. The covariance matrix,
$\mathsf{C}_{ij}=-\frac{\partial\langle q_{j}\rangle_{\underline{\beta}}}{\partial\beta^{i}}=\langle Q_{i},q_{j}\rangle^{\rm c}_{\underline{\beta}}=\int{\rm d}{\rm x}\,\langle q_{i}({\rm x}),q_{j}(0)\rangle_{\underline{\beta}}^{\rm c}=\int{\rm d}{\rm x}\,\big{(}\langle q_{i}({\rm x})q_{j}(0)\rangle_{\underline{\beta}}-\langle q_{i}({\rm x})\rangle_{\underline{\beta}}\langle q_{j}(0)\rangle_{\underline{\beta}}\big{)}$ (10)
is assumed to be positive,
$\mathsf{C}^{\rm T}=\mathsf{C},\quad\mathsf{C}>0,$ (11)
so that the map $\underline{\beta}\mapsto\langle\underline{q}\rangle_{\underline{\beta}}$ is invertible. This defines functions $\beta^{i}(\underline{\mathsf{q}})=\beta^{i}(\mathsf{q}_{1},\mathsf{q}_{2},\ldots)$ for all $i$’s. There is equivalence between microcanonical and macrocanonical ensembles [32, 33, 34]: for every local obsevable $o$, we have $\langle o\rangle_{\underline{\mathsf{q}}}=\langle o\rangle_{\underline{\beta}(\underline{\mathsf{q}})}$. Throughout, we denote the resulting average of $o$, as a function of conserved densities $\underline{\mathsf{q}}$, by using the “sans-serif” font for the observable,
$\mathsf{o}(\underline{\mathsf{q}}):=\langle o\rangle_{\underline{\mathsf{q}}}=\langle o\rangle_{\underline{\beta}(\underline{\mathsf{q}})}.$ (12)
See Appendix A for hydrodynamic matrices, velocities and normal modes.
For our general discussion of hydrodynamics, we consider states which are not space-time stationary, but with variations on long wavelengths $\ell$. It is useful to think of states of the form
$\langle\cdots\rangle_{\ell}=\frac{\int{\rm d}\mu\,{\rm e}^{-\sum_{i}\int{\rm d}{\rm x}\,\beta^{i}({\rm x}/\ell)q_{i}(\bm{{\rm x}})}\cdots}{\int{\rm d}\mu\,{\rm e}^{-\sum_{i}\int{\rm d}{\rm x}\,\beta^{i}({\rm x}/\ell)q_{i}(\bm{{\rm x}})}}$ (13)
for a large, macroscopic scale $\ell$. This state has short range correlations. More general states, with long-range correlations or with long-wavelength source insertions at various macroscopic times, can also be considered.
3 Local relaxation with fluctuations
At the basis of hydrodynamics is the idea of separation of scales: some observables vary slowly in space and time, while other vary much more quickly. By “integrating out” the fast observables, there remains a theory for slow observables, which are the local conserved densities admitted by the system. The Boltzmann Gibbs principle for hydrodynamics at the Euler scale says that linear fluctuations of observables project onto those of densities under such a fast relaxation [6]. Its nonlinear version, dubbed relaxation of fluctuations [25], gives access to large deviations. Beyond the Euler scale, it is well known that one must add fluctuations and diffusive effects. In this section we give heuristic arguments giving such a nonlinear fluctuating Boltzmann-Gibbs principle. We consider the ballistic scaling of space and time, and its corrections in $1/\ell$, but most of this discussion is relevant to other scalings as well, such as the diffusive scaling, see [35, 36, 21, 20, 23].
3.1 Nonlinear fluctuating Boltzmann-Gibbs principle
A mesoscopic region in one dimension is an interval of size $L$ much larger than microscopic lengths $\ell_{\rm micro}$ but much smaller than variation lengths $\ell$,
$\ell\gg L\gg\ell_{\rm micro}.$ (14)
We may consider $\ell_{\rm micro}$ to be finite (say, set to 1). We write
$V_{L}=[-L/2,L/2]$ (15)
and define the fluid-cell mean of an observable – the coarse grained observable – at macroscopic space-time coordinates $x,t$ as
$\overline{o}(x,t):=\frac{1}{L}\int_{V_{L}}{\rm d}{\rm y}\,o(\ell x+{\rm y},\ell t).$ (16)
This precise form of the fluid cell is not mandatory; one may for instance consider a weight function $w({\rm x})$ such that $|w({\rm x})|<a{\rm e}^{-b|x|}$ for some $a,b>0$ and $\int_{\mathbb{R}}{\rm d}{\rm x}\,w({\rm x})=1$, and define
$\overline{o}(x,t):=\frac{1}{L}\int_{{\mathbb{R}}}{\rm d}{\rm y}\,w({\rm y}/L)o(\ell x+{\rm y},\ell t).$ (17)
We argue in Sec. 3.2 that a nonlinear, fluctuating Boltzmann-Gibbs principle holds. That is, coarse-grained observables can be written solely in terms of fields representing coarse-grained conserved densities, which we denote $q_{i}(x,t)$, and emergent Gaussian white noise fields $\xi_{o}(x,t)$’s representing the “forgotten” microscopic degrees of freedom for the observable $o$. Both of these families of fields fluctuate, and we represent averages in the corresponding emergent fluctuating theory as $\langle\!\langle\cdots\rangle\!\rangle_{\ell}$. We propose that one can make the following replacement, which has three parts, microcanonical, diffusive and stochastic (noise):
$\overline{o}(x,t)\ \mbox{in $\langle\cdots\rangle_{\ell}$}$
$\longrightarrow$ $o(x,t):=\underbrace{\mathsf{o}^{\rm reg}(\underline{q}(x,t))}_{\text{microcanonical}}\underbrace{-\frac{\operatorname{sgn}t}{2\ell}\sum_{i}\hat{\mathfrak{D}}_{o}^{~{}i}(\underline{q}(x,t))\partial_{x}q_{i}(x,t)}_{\text{diffusive}}+\underbrace{\frac{1}{\ell}\xi_{o}(x,t)}_{\text{noise}}\ \mbox{in $\langle\!\langle\cdots\rangle\!\rangle_{\ell}$}$
with simply
$\overline{q_{i}}(x,t)\longrightarrow q_{i}(x,t)$ (19)
for conserved densities. Here,
$\langle\!\langle\xi_{o}(x,t)\rangle\!\rangle_{\ell}=0,\quad\langle\!\langle\xi_{o}(x,t)\xi_{o^{\prime}}(x^{\prime},t^{\prime})\rangle\!\rangle_{\ell}=\langle\!\langle\hat{\mathfrak{L}}_{o,o^{\prime}}\big{(}\underline{q}(x,t)\big{)}\rangle\!\rangle\,\delta(x-x^{\prime})\delta(t-t^{\prime})$ (20)
for some noise covariance $\hat{\mathfrak{L}}_{o,o^{\prime}}(\underline{q}(x,t))$, function of the fluctuating densities. $\hat{\mathfrak{D}}_{o}^{~{}i}(\underline{q}(x,t))$ are the “bare” diffusion functions for the observable $o$ which accounts for fluctuation-dissipation effects from the noise and are irreversible (thus the $\operatorname{sgn}(t)$ factor), and $\mathsf{o}^{\rm reg}(\underline{q}(x,t))$ is a regularisation of the microcanonical average (12) in the fluid cell which account for non-fluctuating, non-diffusive corrections due to finite size of the cell. As we will see, the latter is crucial in order to make sense of singularities that arise because of short-range correlations, either as introduced by the initial state, or by the noise. The meaning of the various terms in (3.1) is discussed at more length in Sec. 3.2. These terms will be evaluated in Sec. 4 in terms of equilibrium quantities.
Eq. (3.1) is closely related to nonlinear fluctuating hydrodynamics [9], and a principle of the form (3.1) was proposed in [25, Sec 3.3] as an extension of the Ballistic Macroscopic Fluctuation Theory (BMFT) [25] to include diffusive effects. However in both cases, the constitutive functions $\hat{\mathfrak{L}}_{o,o^{\prime}}\big{(}\underline{q}(x,t)\big{)}$ and $\hat{\mathfrak{D}}_{o}^{~{}i}(\underline{q}(x,t))$ representing the corrections to the Euler scale are not evaluated, and the necessary regularisation $\mathsf{o}^{\rm reg}(\underline{q}(x,t))$ is not discussed. The “bare” noise in (3.1) was also studied in [37]. In systems that are purely diffusive, for current observables $o=j_{k}$ only the diffusive and stochastic part remain, as the microcanonical part is independent of $\underline{q}(x,t)$. This is the basis for the Macroscopic Fluctuation Theory (MFT) [8], a large-deviation theory for fluctuations on diffusive scales.
Recall that $\ell$ is the macroscopic scale: it is present both in $\overline{o}(x,t)$, via the ballistic scaling of space and time, and in the state $\langle\cdots\rangle_{\ell}$, its length scale for variations. The replacement (3.1) gives the $-\ell$ asymptotic expansion of macroscopic physical quantities evolved in time $t$. For instance, the theory gives the asymptotic (distributional) expansion of connected correlation functions: with (3.1) for $o_{k}(x_{k},t_{k})$, we have
$\langle\overline{o_{1}}(x_{1},t_{1}),\ldots,\overline{o_{n}}(x_{n},t_{n})\rangle^{\rm c}_{\ell}=\langle\!\langle o_{1}(x_{1},t_{1}),\ldots,o_{n}(x_{n},t_{n})\rangle\!\rangle^{\rm c}_{\ell}+\mathcal{O}(\ell^{-1-n})$ (21)
with
$\langle\!\langle o_{1}(x_{1},t_{1}),\ldots,o_{n}(x_{n},t_{n})\rangle\!\rangle^{\rm c}_{\ell}=\ell^{1-n}S_{o_{1},\ldots,o_{n}}(x_{1},t_{1};\cdots;x_{n},t_{n})+\ell^{-n}\delta S_{o_{1},\ldots,o_{n}}(x_{1},t_{1};\cdots;x_{n},t_{n})+\ldots$ (22)
See [25, 26] for the general principles of this expansion and the leading Euler-scale scaling. Similarly, for the large deviations and first subleading correction of macroscopic observables such as total currents, we have
$\log\langle{\rm e}^{\lambda\int_{0}^{\ell T}{\rm d}\mathrm{t}\,j_{k}(0,\mathrm{t})}\rangle_{\ell}=\log\langle\!\langle{\rm e}^{\lambda\ell\int_{0}^{T}{\rm d}t\,j_{k}(0,t)}\rangle\!\rangle_{\ell}+\mathcal{O}(\ell^{-1})=\ell F_{k}(\lambda,T)+\delta F_{l}(\lambda,T)+\mathcal{O}(\ell^{-1}).$ (23)
We leave for future work the evaluation of such correlation functions and large-deviation functions, and in particular the corrections to the Euler-scale projection formula developed in [26].
Crucially, the state $\langle\!\langle\cdots\rangle\!\rangle_{\ell}$, which we have not specified except for (20), is in fact completely determined by the initial state correlations $S_{q_{i_{1}},\ldots,q_{i_{n}}}(x_{1},0;\cdots,x_{n},0)$ and $\delta S_{q_{i_{1}},\ldots,q_{i_{n}}}(x_{1},0;\cdots,x_{n},0)$, the replacement (3.1) (where all parts are specified as per Sec 4), and the conservation laws
$\partial_{t}q_{i}(x,t)+\partial_{x}j_{i}(x,t)=0.$ (24)
Indeed, this is because these then provide the dynamics and hence the correlations of conserved densities at later times: this is through the nonlinear fluctuating hydrodynamic equation
$\partial_{t}q_{i}(x,t)+\partial_{x}\Big{(}\mathsf{j}_{i}^{\rm reg}(\underline{q}(x,t))-\frac{\operatorname{sgn}t}{2\ell}\sum_{k}\hat{\mathfrak{D}}_{i}^{~{}k}(\underline{q}(x,t))\partial_{x}q_{k}(x,t)+\frac{1}{\ell}\xi_{j_{i}}(x,t)\Big{)}=0$ (25)
where we use the more standard notation
$\hat{\mathfrak{D}}_{i}^{~{}k}:=\hat{\mathfrak{D}}_{j_{i}}^{~{}k}$ (26)
for the diffusion functions of current observable $j_{i}$.
One way of specifying the state more explicitly is using a path integral formulation such as in the MFT or the BMFT. For a principle of the form (3.1), this was proposed in [25, Sec 3.3], and the path integral formulation was written there, see in particular Eqs. (62)-(65) in that reference. Here, we will not make use of this path integral formulation and the associated large-deviation analysis – it turns out that for all large-deviation results and their subleading corrections, the techniques of cumulant expansion, initially suggested in [25], and fully developed at the Euler scale in [26], is sufficient.
Remark 3.1.
Usually one only considers noise terms associated to the local currents. However, in general, every observable $o$ has its associated, independent noise $\xi_{o}$, generically correlated to all other observables, including currents. It is simple to see from the evolution equation (25) that noise fields for observables that are not currents decouple from the theory, so may be omitted if one looks only at densities and currents.
Remark 3.2.
As usual, although $\xi_{o}$ are white noise fields, the theory is valid at the nonlinear level and for large deviations of local observables. Indeed, white noise fields occur because of the central limit theorem used in the argument of local relaxation to the microcanonical average, but the large-scale fluctuations of conserved densities give rise to nonlinearities and large deviations.
3.2 Coarse graining and the origin of hydrodynamic noise
We now provide general arguments for why (3.1) holds. Although these general argument suggests the form of each correction, they do not allow us to compute them. However, Eq. (3.1), as a principle for the $-\ell$ asymptotic expansion of connected correlation functions, give us a solid framework from which we will evaluate à la Kubo, in the next section, all constitutive elements in terms of equilibrium correlation functions, in linearly degenerate systems.
Consider the fluid-cell mean of a conserved density $\overline{q_{i}}(x,t)$. By the conservation law, its time derivative, with respect to time $\mathrm{t}=\ell t$ at position ${\rm x}=\ell x$ in microscopic units, a is
$\frac{\partial}{\partial\mathrm{t}}\overline{q_{i}}(x,t)=L^{-1}(j_{i}({\rm x}-L/2,\ell t)-j_{i}({\rm x}+L/2,\ell t)).$ (27)
Assuming finite densities and finite typical microscopic velocities, the current observables are of order 1 and vary on a time scale $T$ of order 1 as well. Therefore, $\overline{q_{i}}(x,t)$ has variations of order $\mathcal{O}(L^{-1})$, and this happen on a time scale of order 1. That is, on $T=\mathcal{O}(1)$, $\overline{q_{i}}(x,t)$ can be considered to be invariant, and in particular non-fluctuating.
Now consider the fluid-cell mean of an observable that is not conserved, $\overline{o}(x,t)$. Because it is composed of $\mathcal{O}(L)$ essentially independent observables, each affected by independent fluctuations, the quantity $\overline{o}(x,t)$ varies on a time scale of order $1/L$, each variation being of order $1/L$. Therefore, $\overline{o}(x,t)$ fluctuates quickly, with $\mathcal{O}(L)$ variations, coming from “interaction events” such as collision of particles within the fluid cell, on a time $T$ of order 1. Averaging over $T$, which averages over these fluctuations, the fluid-cell mean of an arbitrary observable should relax to its microcanonical value $\mathsf{o}(\overline{q_{1}}(x,t),\overline{q_{2}}(x,t),\ldots)$.
In fact, by typicality [38, 39, 40] the average over $T$ can be argued not to be necessary. This is because typicality ensures that an average on space implements an average over time in the past, as randomness in the spatial distribution encodes what happened in the past. That is, the $\mathcal{O}(L)$ fluctuations of $\overline{o}(x,t)$ over time $T=\mathcal{O}(1)$ re-organise the system’s configuration in $V_{L}$, under the constraint given by the essentially fixed conserved densities. This re-organisation is enough to implement the small variation of the conserved densities and update $\overline{o}(x,t+T/\ell)$ so that it always take its microcanonical value determined by the $\overline{q_{i}}(x,t)^{\prime}s$.
This suggests that
$\overline{o}(x,t)\stackrel{{\scriptstyle\ell\gg L\gg\ell_{\rm micro}}}{{\to}}\mathsf{o}(\overline{\underline{q}}(x,t))=\mathsf{o}(\overline{q_{1}}(x,t),\overline{q_{2}}(x,t),\ldots).$ (28)
This is the hypothesis of local relaxation of fluctuations [24, 25]. It is a nonlinear Boltzmann-Gibbs principle, where fluctuating local observables keep their full nonlinear dependence on the conserved densities.
The above can only be true for $L$ large enough. The larger the mesoscopic length $L$, the more accurate is its microcanonical relaxation. Yet the cell size cannot be larger than the macroscopic variation scale $\ell$, because the relaxation argument made above requires spatial homogeneity. Further, the larger the time $T$, the more accurate is microcanonical relaxation as well, because more re-organisation has occurred. But $T$ must not bee too large so that conserved densities can be considered fixed and non-fluctuating. Our conjecture is that the optimal choice is obtained for
$LT=\mathcal{O}(\ell).$ (29)
For instance, $L=\ell^{1-\epsilon}$ and $T=\ell^{\epsilon}$ for $\epsilon>0$ small. The time $T$ is not finite, so the asymptotic expansion as $\ell\to\infty$ introduces mistakes of order $\epsilon$. The limit $\epsilon\to 0$ is then taken after the asymptotic expansion as $\ell\to\infty$. Another example is $L=\epsilon\ell$ and $T=1/\epsilon$. Here it is the fluid cell that is too large, introducing mistakes of order $\epsilon$, so that, again, the limit $\epsilon\to 0$ is taken after the asymptotic expansion in $\ell$. This is what is meant by the limit $\ell\gg L\gg\ell_{\rm micro}$.
What are the corrections at order $1/\ell$? The argument above suggests that there are three potential corrections.
Corrections to the microcanonical average. Because the cell size is not infinite, the exact average at fixed conserved densities is not the infinite-volume microcanonical average. Hence on the right-hand side of (28) there is a correction
$L^{-1}\delta\mathsf{o}(\overline{\underline{q}}(x,t))$ (30)
that is
$\mathsf{o}^{\rm reg}(\overline{\underline{q}}(x,t)):=\mathsf{o}(\overline{\underline{q}}(x,t))+L^{-1}\delta\mathsf{o}(\overline{\underline{q}}(x,t)).$ (31)
This is the finite-$L$ microcanonical average up to, including, $L^{-1}$ corrections. We will see in Sec. 4.2 that these corrections are indeed of order $L^{-1}$. The result appears to be non-universal, as, contrary to $\ell$ and $\ell_{\rm micro}$, the fluid cell size $L$ is not physical, but simply an artefact of our construction. However, the emergent fluctuation theory $\langle\!\langle\cdots\rangle\!\rangle_{\ell}$ implicitly depends on $L$, and we will show in Sec. 4.2 how, in linearly degenerate systems, these $L$-dependences effectively cancel out and the result is, in fact, universal.
Diffusion. According to the argument above, there is a delay between the slow variations of the densities and relaxation, so that it is $\overline{o}(x,t+T/\ell)$ that takes the microcanonical value $\mathsf{o}^{\rm reg}(\overline{\underline{q}}(x,t))$. The $\ell^{-1}$ expansion gives an additional term, which is expected to be proportional to the spatial variations $\partial_{{\rm x}}\overline{q}_{i}(x,t)=\ell^{-1}\partial_{x}\overline{q}_{i}(x,t)$. Hence on the right-hand side of (28) there is another correction, for $t>0$:
$-\frac{1}{2}\ell^{-1}\sum_{i}\hat{\mathfrak{D}}_{o}^{~{}i}\big{(}\overline{\underline{q}}(x,t)\big{)}\,\partial_{x}\overline{q_{i}}(x,t)$ (32)
for some set of diffusion functions $\hat{\mathfrak{D}}_{o}^{~{}i}(\overline{\underline{q}}(x,t))$ associated to the observable $o$, where the factor $-\frac{1}{2}$ is conventional. For evolution towards $t<0$, we use PT symmetry, giving the factor $\operatorname{sgn}t$ as both $o$ and $q_{i}$’s are assumed to be PT invariant.
Emergent noise. Microcanonical relaxation, at finite $L$, receives stochastic corrections according to the central limit theorem. This is the result of forgetting the microscopic degrees of freedom when projecting onto the conserved densities. Hence on the right-hand side of (28) there is a noise correction,
$\ell^{-1}\xi_{o}(x,t)$ (33)
with the set of $\xi_{o}(x,t)$’s, for all local observables $o$’s, being correlated centered Gaussian white noises, with (20). The scaling in $\ell^{-1}$ is obtained as follows. The time-averaged fluid cell over $L\times T$ has $LT=\mathcal{O}(\ell)$ independent fluctuations, therefore a variance due to the noise of order $\mathcal{O}(\ell^{-1}$). Let us write a stochastic contribution $\xi_{o}^{\rm micro}(\mathrm{x},\mathrm{t})$ in microscopic coordinates, representing the roughness in both space and time. This can be chosen centered, however it is not a white noise. Take its two-point correlations as $\langle\!\langle\xi_{o}^{\rm micro}(\mathrm{x},\mathrm{t})\xi_{o^{\prime}}^{\rm micro}(\mathrm{x}^{\prime},\mathrm{t}^{\prime})\rangle\!\rangle_{\ell}=\hat{\mathfrak{L}}_{o,o^{\prime}}\delta(\mathrm{x}-\mathrm{x}^{\prime})\delta(\mathrm{t}-\mathrm{t}^{\prime})$. Then, the variance of its total over a fluid cell is
${\rm var}\Big{(}\frac{1}{LT}~\int_{-L/2}^{L/2}{\rm d}{\rm x}\int_{0}^{T}\,{\rm d}\mathrm{t}\,\xi_{o}^{\rm micro}(\mathrm{x},\mathrm{t})\Big{)}=\frac{1}{LT}\hat{\mathfrak{L}}_{o,o^{\prime}}=\ell^{-1}\hat{\mathfrak{L}}_{o,o^{\prime}}.$ (34)
A similar calculation holds for covariances. Transforming to macroscopic coordinates, we obtain (33) with (20), and on that scale the noise becomes Gaussian as higher-point correlations are expected to vanish faster with $\ell^{-1}$.
4 Constitutive relations and the hydrodynamic equation
We now evaluate the beyond-Euler components of the nonlinear fluctuating Boltzmann-Gibbs principle (3.1). For this purpose, we evaluate various equilibrium correlation functions. It turns out that in order to evaluate correlation functions, directly using (3.1), the asymptotic form of correlation functions in (22), and the cumulant expansion principles from Malyshev’s formula (see [26]), are sufficient. In particular, the Euler-scale projection formulas from [25, 26] for 2- and 3-point functions are sufficient for quantities in (3.1) with a factor of $\ell^{-1}$.
The results below hold in $d=1$ linearly degenerate systems [11, 12, 13] (see also [14]). These are systems which possess as many Riemann invariants $n_{I}$ as conserved quantities, that is, the Euler-scale hydrodynamic equation can be written in the form (see Appendix A)
$\partial_{t}n_{I}+\mathsf{v}_{I}\partial_{x}n_{I}=0,$ (35)
and such that the hydrodynamic velocity of mode $I$ does not depend on mode $I$ itself,
$\frac{\partial\mathsf{v}_{I}}{\partial n_{I}}=0\quad\forall\;I.$ (36)
In such systems, no shock appear [15, 16], hence solutions to the Euler equation are expected to be unique at all times. But most importantly for our purposes, in such systems:
1. Correlation functions of conserved densities keep their local-equilibrium form even in non-equilibrium states such as those coming from the long-wavelength initial state (13),
$\langle\!\langle q_{i_{1}}(x_{1},t),q_{i_{2}}(x_{2},t)\rangle\!\rangle_{\ell}^{\rm c}=\mathsf{C}_{i_{1},i_{2}}(x_{1},t)\delta(x_{1}-x_{2})+\mbox{regular}.$ (37)
Here the equilibrium covariance matrix is given in (10), and it is to be evaluated at the state characterised by $\underline{\mathsf{q}}(x_{1},t)=\langle\!\langle\underline{q}(x_{1},t)\rangle\!\rangle_{\ell}$. Eq. (37) was shown to hold in linearly degenerate systems using the BMFT in [25] and [19, Suppl Mat].
2. For every local observable, the three-point couplings, Eq. (106), in space-time stationary, maximal entropy states vanish for modes of equal velocities:
$\langle o_{1}^{-},Q_{I},Q_{I^{\prime}}\rangle^{\rm c}_{\underline{\beta}}=0\quad\mbox{if}\ \mathsf{v}_{I}=\mathsf{v}_{I^{\prime}}.$ (38)
This implies that there is no superdiffusion [9, 10], and was shown from linear degeneracy in [19], Supplementary Material Sec. “Degenerate three-point coupling from linear degeneracy”.
Using these properties of linearly degenerate systems, we will show:
  • I.
    Point-splitting regularisation. The microcanonical regularisation is the point-splitting regularisation introduced in [19],
    $\Big{[}q_{i_{1}}(x_{1},t_{1})\cdots q_{i_{k}}(x_{k},t_{k})\Big{]}^{\rm reg}=\frac{1}{k!}\sum_{\sigma\in S(k)}q_{i_{1}}(x_{1}+\sigma_{i}0^{+},t_{1})\cdots q_{i_{k}}(x_{k}+\sigma_{k}0^{+},t_{k})$ (39)
    where $S(k)$ is the set of permutations of $(1,2,\ldots,k)$, and $0^{+}$ is a fixed, positive infinitesimal number.
  • II.
    Projected Kubo formula. The noise covariance coefficient is given by the projected Kubo formula: with $\langle\cdots\rangle=\langle\cdots\rangle_{\underline{\beta}(\underline{\mathsf{q}})}$ the stationary, maximal entropy, macrocanonical ensemble with averages $\langle q_{i}\rangle=\mathsf{q}_{i}$ of conserved densities,
    $\hat{\mathfrak{L}}_{o_{1},o_{2}}(\underline{\mathsf{q}})=\mathfrak{L}_{o_{1},o_{2}}(\underline{\mathsf{q}})-\frac{1}{2}\sum_{I,I^{\prime}:\mathsf{v}_{I}\neq\mathsf{v}_{I}^{\prime}}\frac{\langle o_{1}^{-},Q_{I},Q_{I^{\prime}}\rangle^{\rm c}\langle o_{2}^{-},Q_{I},Q_{I^{\prime}}\rangle^{\rm c}}{|\mathsf{v}_{I}-\mathsf{v}_{I^{\prime}}|}.$ (40)
    Here $Q_{I}$ are the total conserved quantities in the normal mode basis (see (110)), $\mathsf{v}_{I}$ are the associated hydrodynamic velocities, and $\mathfrak{L}_{o_{1},o_{2}}$ is the conventional Onsager matrix
    $\mathfrak{L}_{o_{1},o_{2}}(\underline{\mathsf{q}}):=\int_{-\infty}^{\infty}{\rm d}\mathrm{t}\,\Big{(}\int{\rm d}{\rm x}\,\langle o_{1}({\rm x},\mathrm{t}),o_{2}(0,0)\rangle^{\rm c}-\mathsf{D}_{o_{1},o_{2}}\Big{)}$ (41)
    with “Drude weight”
    $\mathsf{D}_{o_{1},o_{2}}:=\lim_{\mathrm{t}\to\pm\infty}\int{\rm d}{\rm x}\,\langle o_{1}({\rm x},\mathrm{t}),o_{2}(0,0)\rangle^{\rm c}=\sum_{i_{1},i_{2}}\frac{\partial\mathsf{o}_{1}}{\partial\mathsf{q}_{i_{1}}}\frac{\partial\mathsf{o}_{2}}{\partial\mathsf{q}_{i_{2}}}\mathsf{C}_{i_{1},i_{2}}$ (42)
    where the last equality follows by a general theorem [41].
  • III.
    Extended Einstein relation. The diffusion coefficients $\hat{\mathfrak{D}}_{o}^{~{}k}(\underline{\mathsf{q}})$ are related to the noise covariance by an extended Einstein relation
    $\hat{\mathfrak{L}}_{o,j_{i}}(\underline{\mathsf{q}})=\sum_{k}\hat{\mathfrak{D}}_{o}^{~{}k}(\underline{\mathsf{q}})\mathsf{C}_{ki},\quad\hat{\mathfrak{D}}_{o}^{~{}k}(\underline{\mathsf{q}})=\sum_{i}\hat{\mathfrak{L}}_{o,j_{i}}(\underline{\mathsf{q}})\mathsf{C}^{ik}.$ (43)
It is also instructive to evaluate the Drude weight (42) within our formalism, in order to confirm that it is consistent. This is done in Appendix B.
Our derivation of Point I below gives a full justification of the point-splitting regularisation that was proposed in [19].
Our main result is Point II, the projected Kubo formula. Our proof uses a calculation that follows somewhat that done in [27]: the “diffusion from convection” term is subtracted. The subtraction corresponds to projecting out the quadratic charges forming the “scattering subspace” of the “diffusive space” constructed in [10], $\mathcal{H}_{\rm scat}\subset\mathcal{H}_{\rm dif}$. That is, with the projection $\mathbb{P}_{\rm scat}:\mathcal{H}_{\rm scat}\to\mathcal{H}_{\rm dif}$:
$\hat{\mathfrak{L}}_{o_{1},o_{2}}=\big{(}(1-\mathbb{P}_{\rm scat})o_{1},(1-\mathbb{P}_{\rm scat})o_{2}\big{)}_{\rm dif}.$ (44)
In the construction of [10], the Onsager matrix $\mathfrak{L}_{o_{1},o_{2}}=(o_{1},o_{2})_{\rm dif}$ is written as an inner product in $\mathcal{H}_{\rm dif}$. There is a subspace, $\mathcal{H}_{\rm scat}$, that is spanned by the image of a certain bilinear map on the conserved quantities. This subspace is physically interpreted as describing the aggregated effect of nonlinear two-wave scattering events of hydrodynamic modes emanating from initial fluctuations and transported by the Euler equation. This nonlinear scattering and deterministic transport of fluctuations is the source of long-range correlations, as explained in [24]. As these have $\mathcal{O}(\ell^{-1})$ strength (see (22)), they contribute to $\mathfrak{L}_{o_{1},o_{2}}$ at the same order as the correlation between hydrodynamic noise $\xi_{o_{1}}$ and $\xi_{o_{2}}$. This is why this subspace must be subtracted to obtain the noise covariance.
Point III is expected – for current observables $o=j_{k}$, the diffusion functions $\hat{\mathfrak{D}}_{k}^{~{}i}(\underline{q}(x,t))$ (using the notation (26)) are usually assumed to be related to the noise covariance by Einstein’s relation. However, as far as we are aware, this has not been derived before in the context of (3.1). Further, here the formula is more general because it considers diffusion coefficients for arbitrary observables.
We believe the results of Points II and III still hold beyond linearly degenerate systems, with an appropriate projection including certain unbounded elements of the scattering Hilbert space that makes the resulting noise covariance finite. However we do not know of a derivation yet.
We note that in [23], the same projected Kubo formula (40) is obtained for linearly degenerate systems by using a fluctuation principle of the type (3.1) but under the diffusive scaling, instead of the large-deviation scaling. Importantly, the result is verified by explicit calculations in a stochastic model, thus confirming the projected Kubo formula. The calculation methods are quite different, because of the different scaling.
4.1 The hydrodynamic equations at the diffusive scale
We first write down the diffusive-scale hydrodynamic equation from (3.1), taking into account the results above, Points I-III.
The hydrodynamic equation is obtained from this and the conservation laws, written for the long-wavelength state $\langle\cdots\rangle_{\ell}$ (see (13)), whose corresponding average in the emergent fluctuating theory is denoted $\langle\!\langle\cdots\rangle\!\rangle_{\ell}$:
$\partial_{t}\langle\!\langle q_{i}(x,t)\rangle\!\rangle_{\ell}+\partial_{x}\langle\!\langle j_{i}(x,t)\rangle\!\rangle_{\ell}=0.$ (45)
Inserting (3.1) for the case $o=j_{i}$ is this, using $\langle\!\langle\xi_{o}(x,t)\rangle\!\rangle_{\ell}=0$ and the fact that for the diffusive order $\ell^{-1}$ we may use the leading cumulant-expansion form
$\sum_{i}\langle\!\langle\hat{\mathfrak{D}}_{o}^{~{}i}(\underline{q}(x,t))\partial_{x}q_{i}(x,t)\rangle\!\rangle=\sum_{i}\hat{\mathfrak{D}}_{o}^{~{}i}(\langle\!\langle\underline{q}(x,t)\rangle\!\rangle_{\ell})\partial_{x}\langle\!\langle q_{i}(x,t)\rangle\!\rangle_{\ell}+\mathcal{O}(\ell^{-1})$ (46)
and denoting for ligthness $\mathsf{q}_{i}(x,t)=\langle\!\langle q_{i}(x,t)\rangle\!\rangle_{\ell}$, this gives
$\partial_{t}\mathsf{q}_{i}(x,t)+\partial_{x}\langle\!\langle\mathsf{j}_{i}^{\rm reg}(\underline{q}(x,t))\rangle\!\rangle_{\ell}=\frac{1}{2\ell}\sum_{k}\hat{\mathfrak{D}}_{i}^{~{}k}(\underline{\mathsf{q}}(x,t))\partial_{x}\mathsf{q}_{k}(x,t)$ (47)
where we used the notation (26). The quantity $\langle\!\langle\mathsf{j}_{i}^{\rm reg}(\underline{q}(x,t))\rangle\!\rangle_{\ell}$ was evaluated in [19], and depends on the long-range correlations generated by the long-wavelength state [24], which are at the diffusive order $\ell^{-1}$. These long-range correlations themselves satisfy a set of Euler-scale evolution equation. The result is a set of two combined equations:
$\partial_{t}\mathsf{q}_{i}(x,t)+\sum_{k}\mathsf{A}_{i}^{~{}k}(x,t)\partial_{x}\mathsf{q}_{k}(x,t)$ (48)
$+\,\frac{1}{4\ell}\sum_{r,r^{\prime},s,s^{\prime},\pm}\partial_{x}\Big{(}\langle j_{i}^{-},Q_{r},Q_{s}\rangle^{\rm c}_{\underline{\beta}(x,t)}\mathsf{C}^{rr^{\prime}}(x,t)\mathsf{C}^{ss^{\prime}}(x,t)E_{r^{\prime}s^{\prime}}(x\pm 0^{+},x\mp 0^{+};t)\Big{)}$
$=\ \frac{1}{2\ell}\sum_{k}\hat{\mathfrak{D}}_{i}^{~{}k}(\underline{\mathsf{q}}(x,t))\partial_{x}\mathsf{q}_{k}(x,t)$
with
$\partial_{t}E_{rs}(x,y;t)+\partial_{x}\big{(}\mathsf{A}_{r}^{~{}k}(x,t)E_{ks}(x,y;t)\big{)}+\partial_{y}\big{(}E_{rk}(x,y;t)\mathsf{A}_{s}^{~{}k}(y,t)\big{)}$ (49)
$\qquad\qquad\qquad\qquad=-\delta(x-y)\sum_{kl}\langle j_{k},Q_{r},Q_{s}\rangle^{\rm c}_{\underline{\beta}(x,t)}\mathsf{C}^{kl}(x,t)\partial_{x}\mathsf{q}_{l}(x,t).$
Here $\beta^{i}(x,t):=\beta^{i}(\underline{\mathsf{q}}(x,t))$ (see Eq. (12)) and likewise for the hydrodynamic matrices $\mathsf{A}_{i}^{~{}k}(x,t)$ and $\mathsf{C}_{ik}(x,t)$. In the state (13), the initial conditions are
$\beta^{i}(x,0)=\beta^{i}(x),\quad E_{rs}(x,y;0)=0$ (50)
but these equations are more general, and valid also for initial states that admit non-trivial long-range correlations $E_{rs}(x,y;0)\neq 0$. Eqs. (48), (49) generalise [19, Eqs 9,10], which was for integrable models, to the inclusion of the bare diffusion generically present away from integrability.
4.2 Correction to the microcanonical average and point splitting
Let us consider the microcanonical correction $\delta\mathsf{o}({\underline{q}}(x,t))$, Eq. (30), for one-point averages generically in non-equilibrium states. It be evaluated by studying the average of $o=o(0,0)$ in a stationary, maximal entropy state $\langle\cdots\rangle_{\underline{\beta}}$, and taking into account that it has short-range correlations. Recall that by the equivalence of ensembles,
$\mathsf{o}(\underline{\mathsf{q}})=\langle o\rangle_{\underline{\beta}(\underline{\mathsf{q}})}$ (51)
where the left- (right-)hand side is the microcanonical (macrocanonical) ensemble. Here the Lagrange parameters $\beta^{i}=\beta^{i}(\underline{\mathsf{q}})$ are functions of the average conserved densities defined by $\mathsf{q}_{i}=\langle q_{i}\rangle_{\underline{\beta}(\underline{\mathsf{q}})}=\langle\!\langle q_{i}\rangle\!\rangle_{\underline{\beta}(\underline{\mathsf{q}})}$. Let us denote for lightness of notation $\langle\cdots\rangle=\langle\cdots\rangle_{\underline{\beta}(\underline{\mathsf{q}})}$ and likewise $\langle\!\langle\cdots\rangle\!\rangle=\langle\!\langle\cdots\rangle\!\rangle_{\underline{\beta}(\underline{\mathsf{q}})}$. Then
$\mathsf{o}(\langle\!\langle\underline{q}\rangle\!\rangle)=\mathsf{o}(\underline{\mathsf{q}})=\langle o\rangle=\langle\overline{o}\rangle=\langle\!\langle\mathsf{o}\big{(}\underline{q}\big{)}\rangle\!\rangle+\frac{1}{L}\langle\!\langle\delta\mathsf{o}\big{(}{\underline{q}}\big{)}\rangle\!\rangle$ (52)
[…]
$\langle\!\langle q_{i_{1}}(0,0),\cdots,q_{i_{k}}(0,0)\rangle\!\rangle^{\rm c}$ $=$ $\frac{1}{L^{k}}\int_{(V_{L})^{\times k}}\prod_{i=1}^{k}{\rm d}{\rm x}_{i}\,\langle q_{i_{1}}({\rm x}_{1},0)\cdots q_{i_{k}}({\rm x}_{k},0)\rangle^{\rm c}$ (53)
$=$ $\frac{1}{L^{k-1}}\langle Q_{i_{1}},\ldots,Q_{i_{k-1}},q_{i_{k}}\rangle^{\rm c}$
where corrections (not written) are exponentially small as $L\to\infty$. Therefore we may use a cumulant expansion to get
$\langle\!\langle\mathsf{o}\big{(}\underline{q}\big{)}\rangle\!\rangle$ $=$ $\mathsf{o}\big{(}\underline{\mathsf{q}}\big{)}+\frac{1}{2}\sum_{ij}\frac{\partial^{2}\mathsf{o}(\underline{\mathsf{q}})}{\partial\mathsf{q}_{i}\partial\mathsf{q}_{j}}\langle\!\langle q_{i}(0,0)q_{j}(0,0)\rangle\!\rangle_{\underline{\beta}}^{\rm c}+\mathcal{O}(L^{-2})$ (54)
$=$ $\mathsf{o}\big{(}\underline{\mathsf{q}}\big{)}+\frac{1}{2L}\sum_{ij}\frac{\partial^{2}\mathsf{o}(\underline{\mathsf{q}})}{\partial\mathsf{q}_{i}\partial\mathsf{q}_{j}}\mathsf{C}_{ij}+\mathcal{O}(L^{-2})$
where we used (10). Again by the cumulant expansion we have
$\frac{1}{L}\langle\!\langle\delta\mathsf{o}\big{(}{\underline{q}}\big{)}\rangle\!\rangle=\frac{1}{L}\delta\mathsf{o}\big{(}{\underline{\mathsf{q}}}\big{)}+\mathcal{O}(L^{-2})$ (55)
and hence from (51) we identify
$\delta\mathsf{o}\big{(}{\underline{\mathsf{q}}}\big{)}=-\frac{1}{2}\sum_{ij}\frac{\partial^{2}\mathsf{o}(\underline{\mathsf{q}})}{\partial\mathsf{q}_{i}\partial\mathsf{q}_{j}}\mathsf{C}_{ij}.$ (56)
That is, the correction term simply gets rids of the effects of short-range correlations, the singularities arising at equal positions.
In general evaluating correlation functions using (31) with (56) is rather complicated: we need to keep track of all $1/L$ singularities of $\mathsf{o}(\underline{q}(x,t))$, and all corrections $L^{-1}\delta\mathsf{o}\big{(}{\underline{q}}\big{)}$. These should cancel each other to give something universal (independent of $L$). However, there is a trick that simplifies this procedure and shows universality.
Consider a more general non-equilibrium state $\langle\cdots\rangle_{\ell}$ such as (13), and recall the property (37). Then a similar calculation as (53) and (54) shows that, in every such state, $L^{-1}\delta\mathsf{o}\big{(}{\underline{q}}\big{)}$ cancels the $1/L$ singularities of $\mathsf{o}(\underline{q}(x,t))$. But we can implement this cancellation simply as
$\langle\!\langle\mathsf{o}(\underline{q})\rangle\!\rangle_{\ell}+\frac{1}{L}\langle\!\langle\delta\mathsf{o}(\underline{q})\rangle\!\rangle_{\ell}=\langle\!\langle\mathsf{o}^{\rm reg}(\underline{q})\rangle\!\rangle_{\ell}$ (57)
by defining the regularisation as acting, on monomial of conserved densities at equal space-time points, by avoiding the “fat diagonal” region where coordinates are at microscopic distances from each other:
$\Big{[}\overline{q}_{i_{1}}(x,t)\cdots\overline{q}_{i_{k}}(x,t)\Big{]}^{\rm reg}=\frac{1}{L^{k}}\int_{(V_{L})_{\neq}^{\times k}}\prod_{j=1}^{k}\Big{(}{\rm d}{\rm x}_{j}\,q_{i_{j}}(\ell x+{\rm x}_{i},\ell t)\Big{)}$ (58)
with
$(V_{L})_{\neq}^{\times k}=\{(x_{1},\ldots,x_{k})\in(V_{L})^{\times k}:|x_{a}-x_{b}|\gg\ell_{\rm micro}\,\forall a\neq b\}.$ (59)
Indeed, it is the integration on this fat diagonal that gives the $1/L$ singularity of (53). But, out of equilibrium, inside any sector of a given ordering of spatial coordinates, $x_{\sigma_{1}}>\cdots>x_{\sigma_{k}}$ for some $\sigma\in P(k)$, correlations are continuous at the Euler scale, according to BMFT (see e.g. [17, 19]). As these correlations give $\ell^{-1}$ subleading corrections to the representation of the observable $o$ within the fluctuating theory, they are sufficient to understand the diffusive scale. Therefore, we may replace the regularised fluid-cell mean (58), where the fat diagonal is omitted, simply by the point-splitting regularisation (39).
As we can reproduce correlation functions by functional differentiations with respect to external fields in such states (see [42, 26]), this regularisation works for correlation functions as well. This is a universal procedure: the explicit $L$ dependence has disappeared, and all equal-point singularities are avoided.
Therefore, we have derived the point-splitting regularisation (39) conjectured in [19] from the finite-size correction to the microcanonical averaging in fluid cells.
4.3 Noise correlations via projected Kubo formula
We now evaluate the correlation matrix $\hat{\mathfrak{L}}_{o,o^{\prime}}$ in (20). We will show that, in one dimension, it is not the usual Kubo formula, but instead a projected Kubo formula, where the quadratic charge subspace [10] has been projected out.
[…]
$\mathfrak{L}_{o_{1},o_{2}}$ $=$ $\int_{-\infty}^{\infty}{\rm d}\mathrm{t}\,\Big{(}\int{\rm d}{\rm x}\,\langle o_{1}({\rm x},\mathrm{t}),o_{2}(0,0)\rangle^{\rm c}-\mathsf{D}_{o_{1},o_{2}}\Big{)}$ (60)
$=$ $\ell\int_{-\infty}^{\infty}{\rm d}t\,\Big{(}\ell\int{\rm d}x\,\langle\!\langle o_{1}(x,t),o_{2}(0,0)\rangle\!\rangle^{\rm c}-\mathsf{D}_{o_{1},o_{2}}\Big{)}$
where we have written it in terms of the emergent fluctuating theory, with scaled space-time coordinates, where we will use (3.1). We note that the Drude weight (42) can be written as
$\mathsf{D}_{o_{1},o_{2}}=\ell\sum_{i_{1},i_{2}}\frac{\partial\mathsf{o}_{1}}{\partial\mathsf{q}_{i_{1}}}\frac{\partial\mathsf{o}_{2}}{\partial\mathsf{q}_{i_{2}}}\int{\rm d}x\,\langle\!\langle q_{i_{1}}(x,t),q_{i_{2}}(0,0)\rangle\!\rangle^{\rm c}$ (61)
for every $t\in{\mathbb{R}}$, therefore we may write
$\mathfrak{L}_{o_{1},o_{2}}=\ell^{2}\int_{-\infty}^{\infty}{\rm d}t\int{\rm d}x\,\langle\!\langle\,o_{1}^{-}(x,t),o_{2}^{-}(0,0)\,\rangle\!\rangle^{\rm c}$ (62)
in terms the observable $o$ from which we project out its overlap with the conserved densities:
$o^{-}=o-\sum_{i,j}\langle o,Q_{i}\rangle^{\rm c}\mathsf{C}^{ij}q_{j}=o-\sum_{i}\frac{\partial{\mathsf{o}}}{\partial\mathsf{q}_{i}}q_{i}.$ (63)
Using (3.1), we analyse the terms that contribute to $\langle\!\langle o_{1}(x_{1},t_{1}),o_{2}(x_{2},t_{2})\rangle\!\rangle^{\rm c}$ at orders $\ell^{-1}$ and $\ell^{-2}$; and in particular, those that contribute to (60) (equivalently (62)), hence that are nonzero under space-time integration. For each of the observables $o_{1}$ and $o_{2}$, we consider the microcanonical, diffusive and noise terms, and evaluate their correlations.
First, the microcanonical-microcanonical correlation must be evaluated to leading and first subleading order in $\ell^{-1}$, by the cumulant expansion (or Malyshev’s formula), see the principles explained in [26]. The result is
$\langle\!\langle\mathsf{o}_{1}^{\rm reg}(\underline{q}(x_{1},t_{1})),\mathsf{o}_{2}^{\rm reg}(\underline{q}(x_{2},t_{2}))\rangle\!\rangle^{\rm c}$
$\sim$ $\sum_{i_{1},i_{2}}\frac{\partial\mathsf{o}_{1}}{\partial\mathsf{q}_{i_{1}}}\frac{\partial\mathsf{o}_{2}}{\partial\mathsf{q}_{i_{2}}}\langle\!\langle q_{i_{1}}(x_{1},t_{1}),q_{i_{2}}(x_{2},t_{2})\rangle\!\rangle^{\rm c}$
$+\,\frac{1}{2}\sum_{i_{1},i_{2},i_{1}^{\prime},i_{2}^{\prime}}\frac{\partial^{2}\mathsf{o}_{1}}{\partial\mathsf{q}_{i_{1}}\partial\mathsf{q}_{i_{1}^{\prime}}}\frac{\partial^{2}\mathsf{o}_{2}}{\partial\mathsf{q}_{i_{2}}\partial\mathsf{q}_{i_{2}^{\prime}}}\langle\!\langle q_{i_{1}}(x_{1},t_{1}),q_{i_{2}}(x_{2},t_{2})\rangle\!\rangle^{\rm c}\langle\!\langle q_{i_{1}^{\prime}}(x_{1},t_{1}),q_{i_{2}^{\prime}}(x_{2},t_{2})\rangle\!\rangle^{\rm c}$
$+\,\frac{1}{2}\sum_{i_{1},i_{2},i_{1}^{\prime}}\frac{\partial^{2}\mathsf{o}_{1}}{\partial\mathsf{q}_{i_{1}}\partial\mathsf{q}_{i_{1}^{\prime}}}\frac{\partial\mathsf{o}_{2}}{\partial\mathsf{q}_{i_{2}}}\frac{1}{2}\sum_{\pm}\langle\!\langle q_{i_{1}}(x_{1}\pm 0^{+},t_{1}),q_{i_{1}^{\prime}}(x_{1}\mp 0^{+},t_{1}),q_{i_{2}}(x_{2},t_{2})\rangle\!\rangle^{\rm c}$
$+\,\frac{1}{2}\sum_{i_{1},i_{2},i_{2}^{\prime}}\frac{\partial\mathsf{o}_{1}}{\partial\mathsf{q}_{i_{1}}}\frac{\partial^{2}\mathsf{o}_{2}}{\partial\mathsf{q}_{i_{2}}\partial\mathsf{q}_{i_{2}^{\prime}}}\frac{1}{2}\sum_{\pm}\langle\!\langle q_{i_{1}}(x_{1},t_{1}),q_{i_{2}}(x_{2}\pm 0^{+},t_{2}),q_{i_{2}^{\prime}}(x_{2}\mp 0^{+},t_{2})\rangle\!\rangle^{\rm c}+\mathcal{O}(\ell^{-3})$
where the first line on the right-hand side is $\mathcal{O}(\ell^{-1})$ and the last three lines are $\mathcal{O}(\ell^{-2}$). In the second line, the point-splitting prescription has been dropped, as it will become clear now that it does not influence the result. Expression (4.3) simplifies drastically when written for the projected-out observables (62), and this fully accounts for the Drude weight subtraction:
$\langle\!\langle(\mathsf{o}_{1}^{-})^{\rm reg}(\underline{q}(x_{1},t_{1})),(\mathsf{o}_{2}^{-})^{\rm reg}(\underline{q}(x_{2},t_{2}))\rangle\!\rangle^{\rm c}$
$=$ $\frac{1}{2}\sum_{i_{1},i_{2},i_{1}^{\prime},i_{2}^{\prime}}\frac{\partial^{2}\mathsf{o}_{1}}{\partial\mathsf{q}_{i_{1}}\partial\mathsf{q}_{i_{1}^{\prime}}}\frac{\partial^{2}\mathsf{o}_{2}}{\partial\mathsf{q}_{i_{2}}\partial\mathsf{q}_{i_{2}^{\prime}}}\langle\!\langle q_{i_{1}}(x_{1},t_{1}),q_{i_{2}}(x_{2},t_{2})\rangle\!\rangle^{\rm c}\langle\!\langle q_{i_{1}^{\prime}}(x_{1},t_{1}),q_{i_{2}^{\prime}}(x_{2},t_{2})\rangle\!\rangle^{\rm c}+\mathcal{O}(\ell^{-3}).$
On the right-hand side, we set $x_{1}=x,t_{1}=t$ and $x_{2}=t_{2}=0$, we evaluate
$\frac{\partial^{2}\mathsf{o}_{1}}{\partial\mathsf{q}_{i_{1}}\partial\mathsf{q}_{i_{1}^{\prime}}}=\sum_{j_{1},j_{1}^{\prime}}\langle o_{1}^{-},Q_{j_{1}},Q_{j_{1}^{\prime}}\rangle^{\rm c}\mathsf{C}^{j_{1},i_{1}}\mathsf{C}^{j_{1}^{\prime},i_{1}^{\prime}}$ (66)
and pass to normal modes using (110), to obtain
$\sum_{I_{1},I_{1}^{\prime},I_{2},I_{2}^{\prime}}\langle o_{1}^{-},Q_{I_{1}},Q_{I_{1}^{\prime}}\rangle^{\rm c}\langle o_{2}^{-},Q_{I_{2}},Q_{I_{2}^{\prime}}\rangle^{\rm c}\langle\!\langle q_{I_{1}}(x,t),q_{I_{2}}(0,0)\rangle\!\rangle^{\rm c}\langle\!\langle q_{I_{1}^{\prime}}(x,t),q_{I_{2}^{\prime}}(0,0)\rangle\!\rangle^{\rm c}.$ (67)
This can be evaluated exactly to its leading $\mathcal{O}(\ell^{-2})$ order by using the Euler-scale formula for the two-point functions (113),
$\sum_{I,I^{\prime}}\ell^{-2}\frac{1}{2}\langle o_{1}^{-},Q_{I},Q_{I^{\prime}}\rangle^{\rm c}\langle o_{2}^{-},Q_{I},Q_{I^{\prime}}\rangle^{\rm c}\delta(x-\mathsf{v}_{I}t)\delta(x-\mathsf{v}_{I^{\prime}}t).$ (68)
Therefore, taking into account (38) from linear degeneracy, we obtain, in (62),
$\mathfrak{L}_{o_{1},o_{2}}$ =$\stackrel{{\scriptstyle\text{micro-micro}}}{{=}}$ $\frac{1}{2}\sum_{I,I^{\prime}}\int{\rm d}t\int{\rm d}x\,\langle o_{1}^{-},Q_{I},Q_{I^{\prime}}\rangle^{\rm c}\langle o_{2}^{-},Q_{I},Q_{I^{\prime}}\rangle^{\rm c}\delta(x-\mathsf{v}_{I}t)\delta(x-\mathsf{v}_{I^{\prime}}t)$ (69)
$=$ $\frac{1}{2}\sum_{I,I^{\prime}:\mathsf{v}_{I}\neq\mathsf{v}_{I^{\prime}}}\int{\rm d}t\,\langle o_{1}^{-},Q_{I},Q_{I^{\prime}}\rangle^{\rm c}\langle o_{2}^{-},Q_{I},Q_{I^{\prime}}\rangle^{\rm c}\delta((\mathsf{v}_{I}-\mathsf{v}_{I^{\prime}})t)$
$=$ $\frac{1}{2}\sum_{I,I^{\prime}:\mathsf{v}_{I}\neq\mathsf{v}_{I^{\prime}}}\frac{\langle o_{1}^{-},Q_{I},Q_{I^{\prime}}\rangle^{\rm c}\langle o_{2}^{-},Q_{I},Q_{I^{\prime}}\rangle^{\rm c}}{|\mathsf{v}_{I}-\mathsf{v}_{I^{\prime}}|}.$
Second, the noise-noise correlation is simply
$\langle\!\langle\xi_{o_{1}}(x_{1},t_{1})\xi_{o_{2}}(x_{2},t_{2})\rangle\!\rangle^{\rm c}=\ \ell^{-2}\hat{\mathfrak{L}}_{o_{1},o_{2}}\delta(x_{1}-x_{2})\delta(t_{1}-t_{2})$ (70)
so that
$\mathfrak{L}_{o_{1},o_{2}}\stackrel{{\scriptstyle\text{noise-noise}}}{{=}}\hat{\mathfrak{L}}_{o_{1},o_{2}}.$ (71)
Finally, we show that all remaining terms are of order $\mathcal{O}(\ell^{-3})$ or higher orders in $\ell^{-1}$. Let us use for lightness of notation
$o^{\rm d}(x,t)=-\frac{\operatorname{sgn}t}{2}\sum_{i}\mathfrak{D}_{o}^{~{}i}(\underline{q}(x,t))\partial_{x}q_{i}(x,t).$ (72)
The diffusive-diffusive correlation is
$\langle\!\langle\ell^{-1}o_{1}^{\rm d}(x_{1},t_{1}),\ell^{-1}o_{2}^{\rm d}(x_{2},t_{2})\rangle\!\rangle^{\rm c}=\mathcal{O}(\ell^{-3})$ (73)
because the two-point function has leading behaviour $\mathcal{O}(\ell^{-1})$, Eq. (22). Similarly, the diffusive-noise correlation is $\langle\!\langle\ell^{-1}o_{1}^{\rm d}(x_{1},t_{1}),\ell^{-1}\xi_{o_{2}}(x_{2},t_{2})\rangle\!\rangle^{\rm c}=\mathcal{O}(\ell^{-3})$ because any function of conserved densities evolved in time has a noisy component, coming from integrating the current noise term in the conservation law (25), that is of order $\ell^{-1}$, thus giving the additional factor $\mathcal{O}(\ell^{-1})$.
The diffusive-microcanonical correlation is more subtle. By the cumulant expansion, it is
$\langle\!\langle o_{1}^{\rm reg}(x_{1},t_{1}),o_{2}^{\rm d}(x_{2},t_{2})\rangle\!\rangle^{\rm c}=-\frac{\operatorname{sgn}t_{2}}{2\ell}\sum_{i_{1},i_{2}}\frac{\partial\mathsf{o}_{1}}{\partial\mathsf{q}_{i_{1}}}\mathfrak{D}_{o_{2}}^{~{}i_{2}}(\mathsf{q})\partial_{x_{2}}\langle\!\langle q_{i_{1}}(x_{1},t_{1}),q_{i_{2}}(x_{2},t_{2})\rangle\!\rangle^{\rm c}+\mathcal{O}(\ell^{-3})$ (74)
where we have used the fact that $\langle\!\langle\partial_{x_{2}}q_{i_{2}}(x_{1},t_{1})\rangle\!\rangle=0$ by homogeneity of the state, so that other terms in the cumulant expansion vanish. This is in general of order $\ell^{-2}$. However, it is a total derivative. Because in (60) we integrate over the spatial coordinate, this contribution vanishes in (60). Finally, the microcanonical-noise correlation is, by the cumulant expansion,
$\langle\!\langle o_{1}^{\rm reg}(x_{1},t_{1}),\ell^{-1}\xi_{o_{2}}(x_{2},t_{2})\rangle\!\rangle^{\rm c}=\ell^{-1}\sum_{i}\frac{\partial\mathsf{o}_{1}}{\partial\mathsf{q}_{i}}\langle\!\langle q_{i}(x_{1},t_{1}),\xi_{o_{2}}(x_{2},t_{2})\rangle\!\rangle^{\rm c}+\mathcal{O}(\ell^{-3})$ (75)
which again is nonzero at order $\ell^{-2}$, because of the $\mathcal{O}(\ell^{-1})$ noisy component to $q_{i}(x_{1},t_{1})$ coming from integrating the hydrodynamic equation. However, under spatial integration the result vanishes:
$\langle\!\langle Q_{i},\xi_{o_{2}}(x_{2},t_{2})\rangle\!\rangle^{\rm c}=0$ (76)
because the total charge $Q_{i}=\int{\rm d}x\,q_{i}(x)$ does not have noisy contribution, as the noise in the hydrodynamic equation appears inside a spatial derivative.
Hence, combining (71) with (69), we obtain
$\mathfrak{L}_{o_{1},o_{2}}=\frac{1}{2}\sum_{I,I^{\prime}:\mathsf{v}_{I}\neq\mathsf{v}_{I^{\prime}}}\frac{\langle o_{1}^{-},Q_{I},Q_{I^{\prime}}\rangle^{\rm c}\langle o_{2}^{-},Q_{I},Q_{I^{\prime}}\rangle^{\rm c}}{|\mathsf{v}_{I}-\mathsf{v}_{I^{\prime}}|}+\hat{\mathfrak{L}}_{o_{1},o_{2}}$ (77)
from which we deduce the noise strength
$\hat{\mathfrak{L}}_{o_{1},o_{2}}=\mathfrak{L}_{o_{1},o_{2}}-\frac{1}{2}\sum_{I,I^{\prime}:\mathsf{v}_{I}\neq\mathsf{v}_{I^{\prime}}}\frac{\langle o_{1}^{-},Q_{I},Q_{I^{\prime}}\rangle^{\rm c}\langle o_{2}^{-},Q_{I},Q_{I^{\prime}}\rangle^{\rm c}}{|\mathsf{v}_{I}-\mathsf{v}_{I^{\prime}}|}.$ (78)
This is our main result: in linearly degenerate one-dimensional systems, the noise strength is given by the projected Kubo formula.
4.4 Extended Einstein relation
Finally, we establish the relation between the diffusion function $\hat{\mathfrak{D}}_{o}^{~{}i}(\underline{q})$ in (3.1) (equiv. (3.1)), and the noise covariance $\hat{\mathfrak{L}}_{o,j_{i}}(\underline{q})$ for the observable $o$ and currents $j_{i}$. For this purpose, let us write, using space-time translation invariance of the state and the conservation laws,
$\mathfrak{L}_{o,j_{i}}$ $=$ $\ell^{2}\int_{-\infty}^{\infty}{\rm d}t\int{\rm d}x\,\langle\!\langle o^{-}(x,t),j_{i}(0,0)\rangle\!\rangle^{\rm c}$ (79)
$=$ $\ell^{2}\int_{-\infty}^{\infty}{\rm d}t\int{\rm d}x\,\langle\!\langle o^{-}(0,0),j_{i}(-x,-t)\rangle\!\rangle^{\rm c}$
$=$ $-\ell^{2}\int_{-\infty}^{\infty}{\rm d}t\int{\rm d}x\,x\partial_{x}\langle\!\langle o^{-}(0,0),j_{i}(-x,-t)\rangle\!\rangle^{\rm c}$
$=$ $\ell^{2}\int_{-\infty}^{\infty}{\rm d}t\int{\rm d}x\,x\partial_{t}\langle\!\langle o^{-}(0,0),q_{i}(-x,-t)\rangle\!\rangle^{\rm c}$
$=$ $\ell^{2}\lim_{T\to\infty}\int{\rm d}x\,x\langle\!\langle o^{-}(0,0),\big{(}q_{i}(-x,-T)-q_{i}(-x,T)\big{)}\rangle\!\rangle^{\rm c}$
$=$ $\ell^{2}\lim_{T\to\infty}\int{\rm d}x\,x\langle\!\langle\big{(}o^{-}(x,T)-o^{-}(x,-T)\big{)},q_{i}(0,0)\rangle\!\rangle^{\rm c}.$
We now evaluate this expression by considering in turn the microcanonical, diffusive and noise parts of $o^{-}(x,T)$ and $o^{-}(x,-T)$.
For the microcanonical part, the only term that remains from (4.3) is the third one,
$\mathfrak{L}_{o,j_{i}}\stackrel{{\scriptstyle\text{micro}}}{{=}}\sum_{\pm^{\prime}}\pm^{\prime}\ell^{2}\lim_{T\to\infty}\int{\rm d}x\,x\frac{1}{2}\sum_{k,k^{\prime}}\frac{\partial^{2}\mathsf{o}}{\partial\mathsf{q}_{k}\partial\mathsf{q}_{k^{\prime}}}\frac{1}{2}\sum_{\pm}\langle\!\langle q_{k}(x\pm 0^{+},\pm^{\prime}T),q_{k^{\prime}}(x\mp 0^{+},\pm^{\prime}T),q_{i}(0,0)\rangle\!\rangle^{\rm c}$ (80)
which we can write, by the inverse of the manipulations (79), as
$\frac{\ell^{2}}{4}\int{\rm d}t\int{\rm d}x\,\sum_{k,k^{\prime}}\frac{\partial^{2}\mathsf{o}}{\partial\mathsf{q}_{k}\partial\mathsf{q}_{k^{\prime}}}\sum_{\pm}\langle\!\langle q_{k}(x\pm 0^{+},t),q_{k^{\prime}}(x\mp 0^{+},t),j_{i}(0,0)\rangle\!\rangle^{\rm c}.$ (81)
We now use the nonlinear projection formula for 3-point functions [25, 26]
$\langle\!\langle q_{k}(x\pm 0^{+},t),q_{k^{\prime}}(x\mp 0^{+},t),j_{i}(0,0)\rangle\!\rangle^{\rm c}$ (83)
$=$ $\sum_{l}\frac{\partial\mathsf{j}_{i}}{\partial\mathsf{q}_{l}}\langle\!\langle q_{k}(x\pm 0^{+},t),q_{k^{\prime}}(x\mp 0^{+},t),q_{l}(0,0)\rangle\!\rangle^{\rm c}$
$+\,\sum_{l,l^{\prime}}\frac{\partial^{2}\mathsf{j}_{i}}{\partial\mathsf{q}_{l}\partial\mathsf{q}_{l^{\prime}}}\langle\!\langle q_{k}(x\pm 0^{+},t),q_{l}(0,0)\rangle\!\rangle^{\rm c}\langle\!\langle q_{k^{\prime}}(x\mp 0^{+},t),q_{l^{\prime}}(0,0)\rangle\!\rangle^{\rm c}.$
The first term, when put in (81), gives a total charge $Q_{l}$ and hence the factor
$\frac{\partial}{\partial\beta^{l}}\langle\!\langle q_{k}(\pm 0^{+},t),q_{k^{\prime}}(\mp 0^{+},t)\rangle\!\rangle^{\rm c}=0$ (84)
which vanishes by the short-range correlations of the stationary state, and by the point-splitting. The second term is
$\frac{\ell^{2}}{2}\int{\rm d}t\int{\rm d}x\,\sum_{k,k^{\prime}}\frac{\partial^{2}\mathsf{o}}{\partial\mathsf{q}_{k}\partial\mathsf{q}_{k^{\prime}}}\sum_{l,l^{\prime}}\frac{\partial^{2}\mathsf{j}_{i}}{\partial\mathsf{q}_{l}\partial\mathsf{q}_{l^{\prime}}}\langle\!\langle q_{k}(x,t),q_{l}(0,0)\rangle\!\rangle^{\rm c}\langle\!\langle q_{k^{\prime}}(x,t),q_{l^{\prime}}(0,0)\rangle\!\rangle^{\rm c}$ (85)
which is exactly the term (4.3) leading to the result (69),
$\mathfrak{L}_{o,j_{i}}\stackrel{{\scriptstyle\text{micro}}}{{=}}\frac{1}{2}\sum_{I,I^{\prime}:\mathsf{v}_{I}\neq\mathsf{v}_{I^{\prime}}}\frac{\langle o^{-},Q_{I},Q_{I^{\prime}}\rangle^{\rm c}\langle j_{i}^{-},Q_{I},Q_{I^{\prime}}\rangle^{\rm c}}{|\mathsf{v}_{I}-\mathsf{v}_{I^{\prime}}|}.$ (86)
For the diffusive part, we use (74)
$\mathfrak{L}_{o,j_{i}}\stackrel{{\scriptstyle\text{diffusive}}}{{=}}-\frac{\ell}{2}\sum_{\pm}\pm\lim_{T\to\infty}\int{\rm d}x\,x\operatorname{sgn}(\pm T)\sum_{k}\hat{\mathfrak{D}}_{o}^{~{}k}(\mathsf{q})\partial_{x}\langle\!\langle q_{k}(x,\pm T),q_{i}(0,0)\rangle\!\rangle^{\rm c}$ (87)
giving, by integration by part,
$\mathfrak{L}_{o,j_{i}}\stackrel{{\scriptstyle\text{diffusive}}}{{=}}\ell\lim_{T\to\infty}\int{\rm d}x\,\sum_{k}\hat{\mathfrak{D}}_{o}^{~{}k}(\mathsf{q})\langle\!\langle q_{k}(x,T),q_{i}(0,0)\rangle\!\rangle^{\rm c}=\sum_{k}\hat{\mathfrak{D}}_{o}^{~{}k}(\mathsf{q})\mathsf{C}_{ki}.$ (88)
Finally, the noise part vanishes, $\langle\!\langle\xi_{o}(x,T),q_{i}(0,0)\rangle\!\rangle^{\rm c}=0$, because $q_{i}(0,0)$ does not have any noisy component.
Hence we arrive at
$\mathfrak{L}_{o,j_{i}}=\frac{1}{2}\sum_{I,I^{\prime}:\mathsf{v}_{I}\neq\mathsf{v}_{I^{\prime}}}\frac{\langle o^{-},Q_{I},Q_{I^{\prime}}\rangle^{\rm c}\langle j_{i}^{-},Q_{I},Q_{I^{\prime}}\rangle^{\rm c}}{|\mathsf{v}_{I}-\mathsf{v}_{I^{\prime}}|}+\sum_{k}\hat{\mathfrak{D}}_{o}^{~{}k}(\mathsf{q})\mathsf{C}_{ki}$ (89)
and using (40) established above, we find (43).
5 Absence of noise in integrable systems
It was recently conjectured that the noise in integrable systems must vanish, see [20, 21, 19]. In particular, in [19] a nonlinear Boltzmann-Gibbs principle in the ballistic $1/\ell$ expansion, as in (3.1), is used without noise, in order to evaluate diffusive-order corrections to the hydrodynamic equation, showing agreement with numerical simulation. That is, in intregable models, the contribution to the diffusive scale comes solely from the long-range correlations induced by the large-scale variations of the initial state [24], which are of order $1/\ell$ and affect the (point-splitted) microcanonical average.
In this section we provide a proof that this is the case. First, we show this using previous results on the Onsager matrix, and then we provide a first-principle argument, based on the fluid-cell relaxation picture of Sec. 3.2. We show that in integrable systems, the noise $\xi_{j_{i}}$ for the currents $j_{i}$, as appears at order $\ell^{-1}$ in (3.1), vanishes
$\xi_{j_{i}}=0$ (90)
that is,
$\hat{\mathfrak{L}}_{o,j_{i}}(\underline{\mathsf{q}})=\hat{\mathfrak{L}}_{j_{i},o}(\underline{\mathsf{q}})=0,\quad\hat{\mathfrak{D}}_{o}^{~{}k}(\underline{\mathsf{q}})=0$ (91)
where the second relation follows from (43).
First proof. We show that this is a consequence of our main result (40), along with the following observation.
In the hard rods model, the Onsager matrix for currents $\mathfrak{L}_{j_{i},j_{k}}$, which is in fact an operator as it is taken in the basis of the conserved hard rods velocity $i=v\in{\mathbb{R}}$, was evaluated exactly, see e.g. [28, 4]. Further, in [29, 30] (see also [31]) the Onsager matrix for currents $\mathfrak{L}_{j_{i},j_{k}}$ was evaluated in integrable models, generalising the hard-rod formula. This was done using a variety of techniques, including a direct calculation from form factors in [30]. Then, it was shown in [10] that the general formula in fact takes the form
$\mathfrak{L}_{j_{i},j_{k}}\stackrel{{\scriptstyle\text{integrable models}}}{{=}}\frac{1}{2}\sum_{I\neq I^{\prime}}\frac{\langle j_{i}^{-},Q_{I},Q_{I^{\prime}}\rangle^{\rm c}\langle j_{k}^{-},Q_{I},Q_{I^{\prime}}\rangle^{\rm c}}{|\mathsf{v}_{I}-\mathsf{v}_{I^{\prime}}|}.$ (92)
Therefore, from these results and (40), we conclude that the noise, at least at the leading order, has vanishing correlations
$\hat{\mathfrak{L}}_{j_{i},j_{k}}\stackrel{{\scriptstyle\text{integrable models}}}{{=}}0.$ (93)
By the Cauchy-Schwartz inequality (see [10] for the inner-product structure of the Kubo formula), we have
$(\hat{\mathfrak{L}}_{o,j_{i}})^{2}\leq\hat{\mathfrak{L}}_{o,o}\hat{\mathfrak{L}}_{j_{i},j_{i}}\stackrel{{\scriptstyle\text{integrable models}}}{{=}}0$ (94)
which shows (91).
Second proof. We provide a general first-principle argument based on relaxation of fluid-cell means and fundamental properties of integrable systems.
In many-body integrable models, there exists an infinite number of commuting, independent “higher” time flows, besides the “usual” time evolution. If the integrable system is Hamitonian, then these are generated by infinitely many linearly and functionally independent commuting, extensive Hamiltonians. In translation-invariant systems, these time flows are also translation invariant. That is, we can evolve any local observable not just in time $t=t_{1}$, but also in times $t_{2},t_{3},\ldots$, which we denote by the commuting maps $\tau_{2}^{t_{2}},\tau_{3}^{t_{3}},\ldots$:
$o(x,t_{1},t_{2},t_{3},\ldots)=\tau_{2}^{t_{2}}\tau_{3}^{t_{3}}\cdots o(x,t_{1}).$ (95)
Crucially, because they are commuting flows, each conserved density $q_{i}$ is also a conserved density with respect to all higher times $t_{k}$, with associated currents $j_{ki}$:
$\partial_{t_{k}}q_{i}(x,t_{1},\ldots,t_{n})+\partial_{x}j_{ki}(x,t_{1},\ldots,t_{n})=0.$ (96)
Further, using this, it was shown in [10, App 4] that the Onsager matrix is invariant under these flows,
$\mathfrak{L}_{\tau_{2}^{t_{2}}\tau_{3}^{t_{3}}\cdots j_{i},o}=\mathfrak{L}_{j_{i},o}.$ (97)
It is simple to see that the same holds for the Onsager matrix from which quadratic charges have been projected out, as these are invariant under all higher time flows,
$\hat{\mathfrak{L}}_{\tau_{2}^{t_{2}}\tau_{3}^{t_{3}}\cdots j_{i},o}=\hat{\mathfrak{L}}_{j_{i},o}.$ (98)
Now recall the argument in Sec. 3.2 for the origin of hydrodynamic noise. On a fluid cell of size $L$, after a time $T$, there has been $\mathcal{O}(LT)=\mathcal{O}(\ell)$ fluctuation events for the fluid-cell mean $\overline{o}$ of the observable $o$, while, thanks to the conservation laws, the conserved densities $\overline{q_{i}}$’s are essentially constant. Hence, taking an average over a time $t\in[0,\tau]$ (or, equivalently, not taking the time average, but using typicality) relaxation to the microcanonical average occurs, plus a noise of variance $\propto 1/\ell$ by the central limit theorem.
Consider $o=j_{i}$ to be a current. Let us change this observable by performing an averaging over $n$ of the higher time flows, with higher times in the range $[0,T^{\prime}]$ for some fixed $T^{\prime}>1$ (not scaling with $\ell$),
$[j_{i}]=\frac{1}{{T^{\prime}}^{n}}\int_{0}^{T^{\prime}}{\rm d}t_{2}\cdots\int_{0}^{T^{\prime}}{\rm d}t_{n+1}\prod_{k=2}^{n+1}\tau_{k}^{t_{k}}j_{i}.$ (99)
Taking the time-average over $t_{1}\in[0,T]$, the fluid-cell mean $\overline{[j_{i}]}$ is the average of the original fluid cell mean $\overline{j_{i}}$ over the time-hyperbox $[0,T]\times[0,T^{\prime}]^{n}$ of dimension $n+1$, with $t_{1}\in[0,T]$ and $t_{k}\in[0,T^{\prime}]$ for $k=2,\ldots,n+1$. On this hyperbox, fluid cell means of conserved densities $\overline{q_{i}}$’s are all essentially constant, because of (96). As each time flow is independent, the fluctuations induced, from typical configurations, by each flow are also independent. Thus, on the space-time hyperbox $[0,L]\times[0,T]\times[0,T^{\prime}]^{n}$ the configuration displays a number of independent fluctuation events (particle collisions, etc.) proportional to the volume of this box, hence $\mathcal{O}(LT{T^{\prime}}^{n})$. That is, the fluid cell mean $\overline{[j_{i}]}$ has undergone ${T^{\prime}}^{n}$ times more fluctuations than $\overline{j_{i}}$. Taking $n$ large after the $\ell\to\infty$ asymptotic, we thus find by the central limit theorem that the noise for the observable $[j_{i}]$ has variance
$\langle\!\langle\xi_{[j_{i}]}(x,t)\xi_{[j_{i}]}(x^{\prime},t^{\prime})\rangle\!\rangle\sim c(T^{\prime})\ell^{-1}\delta(x-x^{\prime})\delta(t-t^{\prime}),\quad c(T^{\prime})=\mathcal{O}({T^{\prime}}^{-n})$ (100)
hence
$\hat{\mathfrak{L}}_{[j_{i}],[j_{i}]}=\mathcal{O}({T^{\prime}}^{-n}).$ (101)
By (98), this means
$\hat{\mathfrak{L}}_{j_{i},j_{i}}=\mathcal{O}({T^{\prime}}^{-n}).$ (102)
As $n$ can be taken as large as we wish and $T^{\prime}>1$, this implies $\hat{\mathfrak{L}}_{j_{i},j_{i}}=0$, and by Cauchy Schwartz shows (91).
6 Conclusion
We have developed a hydrodynamic fluctuation theory for linearly degenerate systems in one spatial dimension, up to, including, the diffusive order $1/\ell$ as $\ell\to\infty$ in the ballistic scaling of space and time, ${\rm x}=\ell x,\,\mathrm{t}=\ell t$. The theory is presented as a stochastic PDE in a $1/\ell$ expansion, which is well defined up to the first subleading order. This is based on a nonlinear fluctuating Boltzmann-Gibbs principle, where generic observables are projected onto conserved densities, with the addition of diffusive and noise terms. This nonlinear fluctuating Boltzmann-Gibbs principle is fully justified by heuristic arguments based on the separation of relaxation time-scales between coarse-grained conserved densities and other coarse-grained observables. Relaxation of arbitrary observables to their microcanonical averages is modified by noise via the central limit theorem, by “bare” diffusion via fluctuation-dissipation, and by a non-fluctuating, non-diffusive correction to the microcanonical average, in inverse power of the size of the fluid cell. The noise covariance is shown to be given by a projected Kubo formula, where diffuve-scale effects of long-range correlations [19] are subtracted (projected out, from a Hilbert space perspective [10]). The diffusion is shown to be obtained by Einstein relation from this, and the microcanonical correction is shown to give a point-splitting regularisation [19] that makes the stochastic PDE, with delta-correlations both in noise and initial conditions, well-defined. We show, in two different ways, that the noise vanishes at the diffusive scale in integrable models, proving conjectures made previously [20, 21, 19]. We write down the full hydrodynamic equation for linearly degenerate systems, including effects of long-range correlations [19] and bare diffusion.
It would be interesting to evaluate the first subleading corrections to correlation functions beyond the Euler (large-deviation) scale, and to ballistic large-deviation functions [24, 25]. It would also be interesting to generalise this to higher dimensions (see [26] for the nonlinear projection principles for correlation functions at the leading Euler scale), where long-range correlations do not affect the diffusive scale but may affect higher scales of the hydrodynamic expansion.
It would also be interesting to extend the present theory to higher orders. There, $\mathfrak{L}$ and $\mathfrak{D}$ will need regularisation, which will need to be specified – does point-splitting still work? A natural conjecture is that in integrable systems, currents $j_{i}$ do not show emergent noise at all orders in the hydrodynamic expansion in $1/\ell$,
$\overline{o}(x,t)\ \mbox{in $\langle\cdots\rangle_{\ell}$}\longrightarrow o(x,t):=\underbrace{\mathsf{o}^{\rm reg}(\underline{q}(x,t))}_{\text{microcanonical}}+\,\mathcal{O}(\ell^{-\infty})\ \mbox{in $\langle\!\langle\cdots\rangle\!\rangle_{\ell}$}.$ (103)
This is related to the natural conjecture that semiclassical Bethe particles [43, 44, 45] give a good representation of generalised hydrodynamics at all orders [45, 46]. It would be interesting to prove this within the present formalism.
Finally, going beyond linearly degenerate systems would be very interesting. It is possible that a regularisation similar to that which we introduced in order to make the fluctuating hydrodynamic theory well defined in its asymptotic expansion in $1/\ell$, be related to the regularisation necessary to make the KPZ equation well defined [47]. We also propose that bare diffusion and hydrodynamic noise are given by similar projection principles even more generally, in one-dimensional systems with relevant nonlinearities, in such a way that, for the hydrodynamic noise, the infinite Onsager matrix (infinity being related to hydrodynamic superdiffusion) is projected to a subspace of the diffusive Hilbert space where it is finite.
Acknowledgments. I am grateful to Friedrich Hübner and Tomohiro Sasamoto for discussions, and especially Takato Yoshimura for exchanging ideas while working on the simultaneous paper [23] where related results are obtained. I thank the Institute of Science Tokyo for hospitaly while part of this work was done. This work was supported by EPSRC under grants EP/W010194/1, and EP/Z534304/1 (ERC advanced grant scheme).
Appendix A Hydrodynamic velocities and normal modes
We recall that derivatives with respect to conserved densities are related to covariances in the macrocanonical ensemble. For instance, by the chain rule, we find
$\frac{\partial\mathsf{o}}{\partial\mathsf{q}_{i}}=\sum_{j}\mathsf{C}^{ij}\langle Q_{j},o\rangle_{\underline{\beta}}^{\rm c}$ (104)
where $\mathsf{C}^{ij}$ is the inverse of the covariance matrix $\mathsf{C}_{ij}$, that is $\sum_{j}\mathsf{C}_{ij}\mathsf{C}^{jk}=\delta_{i}^{~{}k}$, and likewise
$\frac{\partial^{2}\mathsf{o}}{\partial\mathsf{q}_{i}\partial\mathsf{q}_{j}}=\sum_{jl}\mathsf{C}^{ij}\mathsf{C}^{kl}\langle Q_{j},Q_{l},o^{-}\rangle_{\underline{\beta}}^{\rm c},\quad o^{-}:=o-\sum_{ij}q_{i}\mathsf{C}^{ij}\langle Q_{j},o\rangle_{\underline{\beta}}^{\rm c}$ (105)
where
$\langle Q_{j},Q_{l},o^{-}\rangle^{\rm c}_{\underline{\beta}}=\int{\rm d}x{\rm d}x^{\prime}\langle q_{j}(x),q_{l}(x^{\prime}),o^{-}(0)\rangle^{\rm c}_{\underline{\beta}}$ (106)
is the “3-point coupling”.
The flux Jacobian ${\mathsf{A}}$ is the matrix obtained by differentiating the fluxes (average currents):
$\frac{\partial{\mathsf{j}}_{i}}{\partial\mathsf{q}_{l}}={\mathsf{A}}_{i}^{~{}l}=\sum_{k}\mathsf{C}^{lk}\langle Q_{k},j_{i}\rangle^{\rm c}_{\underline{\beta}}.$ (107)
It satisfies
${\mathsf{A}}\mathsf{C}=\mathsf{C}{\mathsf{A}}^{\rm T}$ (108)
as shown in various contexts and levels of generality in [48, 49, 9, 50, 30, 51, 14, 52]. This implies that each element of the flux Jacobian vector has real eigenvalues, because the matrix $\sqrt{\mathsf{C}}^{-1}\mathsf{A}\sqrt{\mathsf{C}}$, obtained by a similarity transformation, is real and symmetric. In particular
${\rm spec}\,\mathsf{A}=\{\mathsf{v}_{I}\}_{I}\subseteq{\mathbb{R}}$ (109)
where we use the capital letter $I$ to indicate that it runs over the normal mode basis. The quantities $\mathsf{v}_{I}$ have the interpretation as the hydrodynamic velocities, or generalised sound velocities, for wave propagation. Using the orthogonal matrix $M$ diagonalising the symmetric matrix $\sqrt{\mathsf{C}}^{-1}\mathsf{A}(\hat{\bm{p}})\sqrt{\mathsf{C}}$, the transformation to the normal modes can be chosen as $\mathsf{R}=M\sqrt{\mathsf{C}}^{-1}$:
$q_{I}=\sum_{j}\mathsf{R}_{I}^{~{}j}q_{j},\quad\mathsf{R}\mathsf{A}\mathsf{R}^{-1}={\rm diag}\,(\mathsf{v}_{I}(\hat{\bm{p}}))_{I},\quad\mathsf{R}\mathsf{C}\mathsf{R}^{\rm T}={\bf 1}.$ (110)
The Euler scale hydrodynamic equation at large scales ${\rm x}=\ell x,\,\mathrm{t}=\ell t$, for local maximal entropy states parametrised by $\mathsf{q}_{i}(x,t)$, takes the form
$\partial_{t}\mathsf{q}_{i}(x,t)+\sum_{j}\mathsf{A}_{i}^{~{}j}\partial_{x}\mathsf{j}_{i}(\underline{\mathsf{q}}(x,t))=0.$ (111)
In certain cases (but not all), it can be diagonalised by choosing appropriate, generically nonlinear functions $n_{I}=n_{I}(\underline{\mathsf{q}})$, giving
$\partial_{t}n_{I}+\mathsf{v}_{I}\partial_{x}n_{I}=0.$ (112)
In all cases, at the linearised level the Euler equation can be diagonalised, and gives rise to the following simple Euler-scale correlation functions in the normal mode basis and for coarse-grained observables (see e.g. [14]),
$\langle\overline{q_{I}}(x,t),\overline{q_{I}}(0,0)\rangle_{\underline{\beta}}^{\rm c}\sim\ell^{-1}\delta(x-\mathsf{v}_{I}t).$ (113)
Appendix B Drude weight from nonlinear fluctuating Boltzmann-Gibbs principle
Recall that as a general theorem [41], the Drude weight is given by the projection onto the conserved quantities:
$\mathsf{D}_{o_{1},o_{2}}=\frac{\partial\mathsf{o}_{1}}{\partial\mathsf{q}_{i_{1}}}\frac{\partial\mathsf{o}_{2}}{\partial\mathsf{q}_{i_{2}}}\mathsf{C}_{i_{1},i_{2}}.$ (114)
It is instructive to verify that our formalism is in agreement with the general result (114), by evaluating the Drude weight using the nonlinear fluctuating Boltzmann-Gibbs principle (3.1).
Consider the definition (42) written in terms of macrosocpic coordinates for fluid-cell means in the fluctuation theory,
$\mathsf{D}_{o_{1},o_{2}}=\lim_{t\to\pm\infty}\ell\int{\rm d}x\,\langle\!\langle o_{1}(x,t),o_{2}(0,0)\rangle\!\rangle^{\rm c}.$ (115)
Consider $\langle\!\langle o_{1}(x_{1},t_{1}),o_{2}(x_{2},t_{2})\rangle\!\rangle^{\rm c}$, which depends on $x_{1}-x_{2}$ and $t_{1}-t_{2}$ by space-time translation invariance. We may integrate over $x_{1}$, or integrate over $x_{2}$, and we look at the limit $t_{1}-t_{2}\to\pm\infty$. Clearly, the noise-boise term (70) gives zero under the limit $t_{1}-t_{2}\to\pm\infty$, and by the results of Sec. 4.3 all other correlations vanish except for that from the microcanonical terms in (3.1), that is (4.3).
For the fist line on the right-hand side of (4.3), we obtain (114). Hence, we need to verify that the other lines give zero.
For the last line on the right-hand side of (4.3), we integrate over $x_{1}$ and notice that the result is independent of $t_{1},t_{2}$:
$\int{\rm d}x_{1}\,\langle\!\langle q_{i_{1}}(x_{1},t_{1}),q_{i_{2}}(x_{2}\pm 0^{+},t_{2}),q_{i_{2}^{\prime}}(x_{2}\mp 0^{+},t_{2})\rangle\!\rangle^{\rm c}$
$=$ $\langle\!\langle Q_{1},q_{i_{2}}(x_{2}\pm 0^{+},t_{2}),q_{i_{2}^{\prime}}(x_{2}\mp 0^{+},t_{2})\rangle\!\rangle^{\rm c}=-\frac{\partial}{\partial\beta^{1}}\langle\!\langle q_{i_{2}}(\pm 0^{+},0),q_{i_{2}^{\prime}}(\mp 0^{+},0)\rangle\!\rangle^{\rm c}=0$
where we use the fact that the state has short-range correlations. Similarly, we obtain 0 for the third line on the right-hand side of (4.3) by integrating over $x_{2}$.
For the second line on the right-hand side of (4.3), we set $x_{1}=x,t_{1}=t$ and $x_{2}=t_{2}=0$, and, as in Sec. 4.3, we pass to normal modes using
$\frac{\partial^{2}\mathsf{o}_{1}}{\partial\mathsf{q}_{i_{1}}\partial\mathsf{q}_{i_{1}^{\prime}}}=\langle o_{1}^{-},Q_{j_{1}},Q_{j_{1}^{\prime}}\rangle^{\rm c}\mathsf{C}^{j_{1},i_{1}}\mathsf{C}^{j_{1}^{\prime},i_{1}^{\prime}}$ (117)
and the normal-mode transformation (110), to obtain
$\langle o_{1}^{-},Q_{I_{1}},Q_{I_{1}^{\prime}}\rangle^{\rm c}\langle o_{2}^{-},Q_{I_{2}},Q_{I_{2}^{\prime}}\rangle^{\rm c}\langle\!\langle q_{I_{1}}(x,t),q_{I_{2}}(0,0)\rangle\!\rangle^{\rm c}\langle\!\langle q_{I_{1}^{\prime}}(x,t),q_{I_{2}^{\prime}}(0,0)\rangle\!\rangle^{\rm c}.$ (118)
This can be evaluated exactly to its leading $\mathcal{O}(\ell^{-2})$ order by using the Euler-scale formula for the two-point functions
$\ell^{-2}\frac{1}{2}\langle o_{1}^{-},Q_{I},Q_{I^{\prime}}\rangle^{\rm c}\langle o_{2}^{-},Q_{I},Q_{I^{\prime}}\rangle^{\rm c}\delta(x-\mathsf{v}_{I}t)\delta(x-\mathsf{v}_{I^{\prime}}t).$ (119)
where we used (63) for the projected observable $o^{-}$. Clearly, the integral over $x$ is 0 if for every values of the indices $I,I^{\prime}$ we have either $\mathsf{v}_{I}\neq\mathsf{v}_{I^{\prime}}$, or $\langle o_{1}^{-},Q_{I},Q_{I^{\prime}}\rangle^{\rm c}=0$, or $\langle o_{2}^{-},Q_{I},Q_{I^{\prime}}\rangle^{\rm c}$. As we assume linear degeneracy, (38) holds, hence the result vanishes.